### v1

# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

Abstract.
Trajectory collection is fundamental for location-based services but often involves sensitive information, such as a user’s daily routine, raising privacy concerns. Local differential privacy (LDP) provides provable privacy guarantees for users, even when the data collector is untrusted. Existing trajectory collection methods ensure LDP only for discrete location spaces, where the number of locations affects their privacy guarantees and trajectory utility. Moreover, the location space is often naturally continuous, such as in flying and sailing trajectories, making these methods unsuitable. This paper proposes two trajectory collection methods that ensure LDP for continuous spaces: TraCS-D, which perturbs the direction and distance of locations, and TraCS-C, which perturbs the Cartesian coordinates of locations. Both methods are theoretically and experimentally analyzed for trajectory utility. TraCS can also be applied to discrete spaces by rounding perturbed locations to the nearest discrete points. It is independent of the number of locations and has only $\Theta(1)$ time complexity in each perturbation generation. Evaluation results on discrete location spaces validate this advantage and show that TraCS outperforms state-of-the-art methods with improved trajectory utility, especially for large privacy parameters.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ZhengYeah/TraCS.
1. Introduction
A trajectory is a sequence of locations that describes movement over time. Users’ trajectory data serves as a fundamental data source for location-based applications, such as navigation and route planning. However, the direct collection of trajectory data from users raises significant privacy concerns, particularly given the growing privacy expectations of users and stringent regulations (gdp, 2016; Krumm, 2009). Anonymization techniques, such as $k$-anonymity (Sweeney, 2002; Li et al., 2007), have been shown to be insufficient for protecting trajectory privacy (Chen et al., 2011). Therefore, stronger privacy guarantees are needed for such data.
Local differential privacy (LDP) provides a provable worst-case privacy guarantee by requiring the hardness of distinguishing the presence of a user’s data from any released data. The privacy level of an LDP mechanism is quantified by its privacy parameter $\varepsilon$, which also implies the extent of data perturbation. Through LDP, users can release their perturbed trajectory data to an untrusted data collector while still benefiting from the location-based services provided by the collector.
Existing LDP methods for trajectory collection are designed for discrete spaces. They either rely on properly splitting the entire location space into grids (Du et al., 2023) or assume the location space consists of a set of labeled locations (a.k.a. points of interest) (Cunningham et al., 2021b; Zhang et al., 2023). These methods’ privacy guarantees are w.r.t. their corresponding discretized sets, and their efficacy depends on the size of these sets. Improper discretization leads to a loss of privacy guarantees and reduced trajectory utility. Moreover, labeled locations are not always available, especially in continuous spaces such as those encountered in flying and sailing trajectories. Therefore, it is necessary to design effective trajectory collection methods for continuous spaces.
However, designing such methods is challenging due to the following reasons: (i) they must guarantee LDP w.r.t. each location in the continuous space; (ii) despite the rigorous privacy guarantee, they must still provide sufficient trajectory utility.
This paper proposes two LDP methods for trajectory collection in continuous spaces. The main insight is to decompose the $2$D continuous space into $1$D subspaces and design mechanisms for each subspace. Our mechanisms leverage the concept of piecewise-based mechanisms (Wang et al., 2019; Li et al., 2020) for $1$D bounded numerical domains.  Based on two different decompositions of the continuous space, we propose two methods, TraCS-D and TraCS-C. TraCS-D treats the continuous space as the composition of a direction space and a distance space. We design a novel direction perturbation mechanism to ensure LDP for the direction space. Combined with a redesigned piecewise-based mechanism on the distance space, TraCS-D ensures LDP for the continuous $2$D space. TraCS-C treats the continuous space as the Cartesian product of two distance spaces. Each location has a $2$D Cartesian coordinate representation, and TraCS-C perturbs these coordinates.  Our methods can also be applied to discrete spaces by rounding the perturbed location to the nearest discrete points. Compared with existing methods based on the Exponential mechanism (Cunningham et al., 2021b; Zhang et al., 2023), our methods are not affected by the number of locations in the location space and cost only $\Theta(1)$ time in each perturbation generation. Specifically, our contributions are as follows:
  • To the best of our knowledge, this is the first work to provide trajectory collection methods for continuous space under LDP. We propose TraCS-D and TraCS-C for this purpose.  These methods introduce novel location perturbation mechanisms that exploit the direction information and coordinate information of trajectories in continuous spaces. We theoretically and experimentally analyze their trajectory utility.
  • Our methods TraCS can also be applied to discrete spaces. Compared with existing methods for discrete spaces, TraCS have significantly lower computation complexity in generating perturbed locations. Experimental results on the discrete location spaces validate the efficiency of TraCS, and also show that it outperforms the state-of-the-art methods with enhanced trajectory utility, particularly for large privacy parameters.
2. Preliminaries
This section formulates the problem, introduces LDP, and reviews existing trajectory collection methods.
2.1. Problem Formulation
A typical trajectory collection schema consists of a set of trajectories from users and one collector. Each trajectory $\mathcal{T}$ is a sequence of locations $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, where $\tau_{i}\in\mathcal{S}$ and $\mathcal{S}\subset\mathbb{R}^{2}$ is a continuous and bounded domain. The collector needs to collect these trajectories to provide location-based services.
However, the collector is untrusted and may act as an adversary attempting to infer users’ private data. Therefore, releasing the private trajectories to the collector pose a privacy risk. To protect privacy, users perturb their trajectories using a privacy mechanism $\mathcal{M}:\mathcal{S}\to Range(\mathcal{M})$, and then send the perturbed trajectories $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$ to the collector.
We aim to design a mechanism $\mathcal{M}$ that satisfies $\varepsilon$-LDP, providing a provable privacy guarantee for the private trajectories. Additionally, $\mathcal{M}$ should ensure the utility of the perturbed trajectories and incur limited computational cost for real-time applications.
2.2. Local Differential Privacy
Definition 2.1 ($\varepsilon$-LDP (Duchi et al., 2013)).
A perturbation mechanism $\mathcal{M}:\mathcal{X}\to Range(\mathcal{M})$ satisfies $\varepsilon$-LDP, iff for two arbitrary input $x_{1}$ and $x_{2}$, the probability ratio of outputting the same $y$ is bounded:
$\forall x_{1},x_{2}\in\mathcal{X},\forall y\in Range(\mathcal{M}):\frac{\Pr[\mathcal{M}(x_{1})=y]}{\Pr[\mathcal{M}(x_{2})=y]}\leq\exp(\varepsilon).$
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function ($pdf$). Intuitively, Definition 2.1 represents the difficulty of distinguishing $x_{1}$ and $x_{2}$ given $y$. Lower values of the privacy parameter $\varepsilon\in[0,+\infty)$ mean higher privacy. For example, $\varepsilon=0$ requires $\mathcal{M}$ to map two arbitrary inputs to any output $y$ with the same probability, thus the perturbed output contains no distribution information of the private input, making any hypothesis-testing method to infer the private $x$ powerless.
Theorem 2.2 (Sequential Composition (Dwork and Roth, 2014; Zhang et al., 2023)).
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two mechanisms that satisfy $\varepsilon_{1}$ and $\varepsilon_{2}$-LDP, respectively. Their composition, defined as $\mathcal{M}_{1,2}:(\mathcal{X}_{1},\mathcal{X}_{2})\to(Range(\mathcal{M}_{1}),Range(\mathcal{M}_{2}))$, satisfies $(\varepsilon_{1}+\varepsilon_{2})$-LDP.
Sequential composition allows us to design LDP mechanisms for $2$D data by composing two mechanisms for $1$D.
2.2.1. Piecewise-based Mechanism
If $\mathcal{X}$ is continuous and bounded, the state-of-the-art LDP mechanisms are piecewise-based (Wang et al., 2019; Li et al., 2020), i.e. they sample the output from piecewise probability distributions defined on $\mathcal{X}$. We summarize them in the following form:
Definition 2.3 (Piecewise-based mechanism).
A piecewise-based mechanism $\mathcal{M}:\mathcal{X}\to Range(\mathcal{M})$ is a family of probability distributions that, given input $x\in\mathcal{X}$, outputs $y\in Range(\mathcal{M})$ according to
$pdf[\mathcal{M}(x)=y]=\begin{dcases}p_{\varepsilon}&\text{if}\ y\in[l_{x,\varepsilon},r_{x,\varepsilon}),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}$ is the sampling probability w.r.t. $\varepsilon$ and $[l_{x,\varepsilon},r_{x,\varepsilon})$ is the sampling interval w.r.t. $x$ and $\varepsilon$. $Range(\mathcal{M})\supseteq\mathcal{X}$ is also a continuous and bounded domain.
Piecewise-based mechanisms sample the output $y$ with a high probability $p_{\varepsilon}$ within $[l_{x,\varepsilon},r_{x,\varepsilon})$ and a lower probability in the remaining two pieces, while satisfying the $\varepsilon$-LDP constraint. Typical instantiations include PM (Wang et al., 2019) for mean estimation and SW (Li et al., 2020) for distribution estimation. Our perturbation mechanisms also leverage the piecewise-based design but with different instantiations tailed for different data domains.
2.2.2. $k$-RR
If $\mathcal{X}$ is discrete, especially when $|\mathcal{X}|$ is not large (Wang et al., 2017), a classical LDP mechanism is $k$-RR (Kairouz et al., 2014).
Definition 2.4.
$k$-RR is a sampling mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{X}$ that, given input $x\in\mathcal{X}$, outputs $y\in\mathcal{X}$ according to
$\Pr[\mathcal{M}(x)=y]=\begin{dcases}\frac{\exp(\varepsilon)}{|\mathcal{X}|-1+\exp(\varepsilon)}&\text{if}\ y=x,\\ \frac{1}{|\mathcal{X}|-1+\exp(\varepsilon)}&\text{otherwise}.\end{dcases}$
$k$-RR outputs the truth $x$ with a higher probability or outputs other values with a lower probability, while satisfying $\varepsilon$-LDP.
2.3. Existing Methods
Table 1. Comparison with existing methods.
Main technique LDP guarantee
NGram (Cunningham et al., 2021b) Hierarchical decomposition Discrete space
LDPTrace (Du et al., 2023) Model-based synthesis
ATP (Zhang et al., 2023) Direction perturbation
This paper (TraCS) Direction* & coordinate perturbation Continuous space
  • * designed for continuous direction space (different from ATP).
Table 1 summarizes existing LDP methods for trajectory collection and our methods.
[…]
LDPTrace (Du et al., 2023) synthesizes trajectories based on a generative model. Specifically, the physical space $\mathcal{S}$ is discretized into cells, and each private trajectory is represented as a sequence of cells $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$. Neighboring cells, i.e. $c_{i}$ and $c_{i+1}$, form transition patterns, which are perturbed by LDP mechanisms. These perturbed transition patterns are then used to train a Markov chain model that captures the transition patterns between cells. Finally, new trajectories $\mathcal{T}^{\prime}$ are synthesized by the trained model.
ATP (Zhang et al., 2023) uses direction perturbation to limit the set of possible locations. It assumes the location space is a finite set of labeled locations and represents a trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$ as a sequence of these locations. When perturbing a location $\tau_{i+1}$, ATP perturbs the direction from $\tau_{i}$ to $\tau_{i+1}$. It divides the direction space into $k$ sectors and uses $k$-RR to sample a direction sector. The perturbed location $\tau_{i+1}^{\prime}$ should be within the sampled sector, thereby limiting the possible $\tau_{i+1}^{\prime}$, which is then sampled using the Exponential mechanism.
We summarize their limitations in the context of continuous spaces as follows:
  • These methods are limited to handling discrete location spaces. The number of locations or cells is a crucial parameter that determines the perturbation mechanism and affects the privacy guarantee. Although it is possible to discretize continuous spaces, proper discretization is challenging to balance the trajectory utility and privacy.
  • Besides the discretization, methods based on the Exponential mechanism, such as NGram and ATP, suffer from high complexity in perturbation generation and trajectory utility loss when the discrete space is large. Additionally, designing public knowledge in NGram is challenging, especially for flying or sailing trajectories; synthesis methods in LDPTrace rely on the regularity of the training trajectories, which is not always guaranteed.
Appendix B.1 details the weaknesses of the Exponential mechanism. In summary, it requires calculating the score functions of each possible pair of locations. This costs $\Theta(m)$ time, where $m$ is the number of locations in the discrete space. Additionally, each sampling (perturbation generation) also costs $\Theta(m)$ time. This time complexity is not affordable for large-scale location spaces.
To address these limitations, we propose TraCS, a novel trajectory collection method that provides LDP guarantees for trajectories in continuous spaces.
3. Our Method: TraCS
This section presents our methods, TraCS-D and TraCS-C, along with their theoretical analyses and extensions.
3.1. Location Space
We consider the location space constrained by a pair of longitude and latitude lines, i.e. $\mathcal{S}\subset\mathbb{R}^{2}$ is a rectangular area $[a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, where $a_{\text{sta}},a_{\text{end}}$ denote the longitudes and $b_{\text{sta}},b_{\text{sta}}$ denote the latitudes. Consequently, each location $\tau_{i}\in\mathcal{S}$ has a natural representation as a pair of coordinates $(a_{i},b_{i})$.  This representation aligns with real-world location data (e.g. GPS data) and existing trajectory collection methods for discrete spaces (Cunningham et al., 2021b; Du et al., 2023; Zhang et al., 2023; Jiang et al., 2022).
The core idea behind having various LDP methods for the location space $\mathcal{S}$ stems from its multiple decompositions. One decomposition means $\mathcal{S}$ can be represented by several subspaces. For example, besides the aforementioned longitude-latitude coordinate representation, each location $\tau_{i}\in\mathcal{S}$ can also be represented by a direction component and a distance component. Based on these decompositions, we propose two LDP methods: TraCS-D, which perturbs the direction and distance subspaces, and TraCS-C, which perturbs subspaces of the Cartesian coordinates.
3.2. TraCS-D
Continuous space $\mathcal{S}$ can be represented as the composition of a direction space $\mathcal{D}_{\varphi}$ and a distance space $\mathcal{D}_{r(\varphi)}$ at each (reference) location $\tau_{i}$. Figure 1(a) illustrates this decomposition.
(a) TraCS-D: $\mathcal{S}=\mathcal{D}_{\varphi}\otimes\mathcal{D}_{r(\varphi)}$
(b) TraCS-C: $\mathcal{S}=\mathcal{D}_{a}\times\mathcal{D}_{b}$
Figure 1. Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$.
[…]
Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$.
Direction space $\mathcal{D}_{\varphi}$. In the space $\mathcal{S}\subset$ $\mathbb{R}^{2}$, the direction is represented by the angle $\varphi$ relative to a reference direction. We define the reference direction as the latitude lines, i.e. the $0^{\circ}$ direction. Thus, the direction space $\mathcal{D}_{\varphi}$ is a circular domain $[0,2\pi)$ for any location. Fixing a location $\tau_{i}$, any other location $\tau_{j}\neq\tau_{i}$ has a unique direction with respect to $\tau_{i}$.
Distance space $\mathcal{D}_{r(\varphi)}$. The other subspace, i.e. the distance space $\mathcal{D}_{r(\varphi)}$, is determined by the direction $\varphi$. Its size is the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ in the direction $\varphi$, which is a function of $\varphi$. For the rectangular $\mathcal{S}$, $r(\varphi)$ has a closed-form expression, which is detailed later. Thus, fixing $\tau_{i}$ and a direction $\varphi$, any other location $\tau_{j}$ has a unique distance $r(\varphi)$ with respect to $\tau_{i}$.
By this decomposition, we can represent any location $\tau_{j}\in S$ as coordinates $(\varphi,r(\varphi))\in(\mathcal{D}_{\varphi},\mathcal{D}_{r(\varphi)})$ relative to $\tau_{i}$. As an example, $\tau_{j}$ in Figure 1(a) is represented by $(\varphi_{1},r(\varphi_{1}))$. This representation is unique and reversible: any pair of coordinates $(\varphi,r(\varphi))$ can be mapped back to a unique location $\tau_{j}\in\mathcal{S}$. Thus, if $\tau_{j}$ is a private location, an LDP mechanism applied to $\tau_{j}=(\varphi,r(\varphi))$ is essentially applied to $\varphi$ and $r(\varphi)$. By Composition Theorem 2.2, we can design perturbation mechanisms for $\varphi$ and $r(\varphi)$, respectively, to achieve LDP for $\tau_{j}$.
(a) $\varepsilon=4$
(b) $\varepsilon=6$
Figure 2. Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$. The red angular arcs indicate the dominant sectors, with their probabilities shown on the right.
[…]
Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$.
3.2.1. Direction Perturbation
The technical challenge in direction perturbation is designing an LDP mechanism for the circular domain $\mathcal{D}_{\varphi}=[0,2\pi)$. This is non-trivial due to the following properties of the circular domain:
  • The circular domain $[0,2\pi)$ is bounded but not finite, making the unbounded mechanisms and discrete mechanisms inapplicable.
  • The circular domain has a unique distance metric, e.g. the distance between 00 and $2\pi$ is 00, not $2\pi$, making the Euclidean distance-based mechanisms inapplicable.
Specifically, classical mechanisms like Laplace and Gaussian (Dwork and Roth, 2014) add unbounded noise, resulting in an output domain of (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). This output domain makes them inapplicable to the circular domain $[0,2\pi)$. Additionally, these mechanisms are designed for the distance metric $d(y,x)=|y-x|$, which is inconsistent with the distance in circular domain. Discrete mechanisms, such as the $k$-RR mechanism (Kairouz et al., 2014), require a finite domain size $|k|$, making them unsuitable for direct application to $[0,2\pi)$.
To illustrate the necessity and superiority of TraCS-D’s piecewise-based design, we start with a strawman approach that extends the $k$-RR mechanism to the circular domain $[0,2\pi)$.
Strawman approach: $k$-RR + uniform sampling. ATP (Zhang et al., 2023) divides the direction space $[0,2\pi)$ into $k$ sectors and applies the $k$-RR mechanism. This approach guarantees LDP for the $k$ sectors but not for the entire $[0,2\pi)$. Even so, we can extend this approach to $[0,2\pi)$ by further uniformly sampling a direction $\varphi^{\prime}$ from the output sector of $k$-RR.
Specifically, the circular domain $[0,2\pi)$ is divided into $k$ sectors $[i-1,i)\cdot 2\pi/k$ for $i=1,2,\ldots,k$. Assume the private direction $\varphi$ falls into the $j$-th sector. Applying the $k$-RR mechanism outputs sector $i=j$ with probability $p=\exp(\varepsilon)/(k-1+\varepsilon)$, or outputs sector $i\neq j$ with probability $p/\exp(\varepsilon)$. Then, we uniformly sample a direction $\varphi^{\prime}$ from the $i$-th sector, i.e. $\varphi^{\prime}\sim U(i-1,i)\cdot 2\pi/k$.
Limitation. This strawman approach introduces inherent inner-sector errors due to uniform sampling within each pre-defined sector. If $k$ is small, uniform sampling within each sector leads to significant inner-sector errors, even for large privacy parameter $\varepsilon$. Conversely, too large $k$ values impair the utility of the $k$-RR mechanism (Wang et al., 2017).
Design rationale. To address this limitation, we design a direction perturbation mechanism in which the perturbed direction is independent of hyperparameters like $k$, thereby avoiding inherent errors beyond those introduced by the privacy parameter $\varepsilon$.
We leverage the idea of piecewise-based mechanisms’ design. Specifically, (i) we instantiate a piecewise-based mechanism for the circular domain $[0,2\pi)$, which guarantees LDP for the entire direction space; (ii) the perturbed direction is sampled from a piecewise probability distribution, centered around the private direction with high probability and only dependent on the privacy parameter $\varepsilon$.
Definition 3.1 (Direction Perturbation Mechanism).
Given a private direction $\varphi$ and a privacy parameter $\varepsilon$, TraCS-D’s direction perturbation mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ is defined by:
$pdf[\mathcal{M}_{\circ}(\varphi)=\varphi^{\prime}]=\begin{dcases}p_{\varepsilon}&\text{if}\ \varphi^{\prime}\in[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon}),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=\frac{1}{2\pi}\exp(\varepsilon/2)$ is the sampling probability, and $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
$\begin{split}l_{\varphi,\varepsilon}&=\left(\varphi-\pi\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}\right)\mod 2\pi,\\ r_{\varphi,\varepsilon}&=\left(\varphi+\pi\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}\right)\mod 2\pi.\end{split}$
The above mechanism $\mathcal{M}_{\circ}$ is defined by a piecewise probability distribution with three pieces, where the central piece $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ has a higher probability density $p_{\varepsilon}$. Obviously, this mechanism satisfies LDP for the circular domain $[0,2\pi)$.
A key advantage of TraCS-D’s direction perturbation mechanism is the “dominant” sector. We refer to the central piece $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ as the dominant sector because it centers around the private direction $\varphi$ with high probability. The dominant sector varies with the privacy parameter $\varepsilon$; a larger $\varepsilon$ results in a narrower dominant sector, closer to the private direction $\varphi$, eliminating the inner-sector errors of the strawman approach.
Example 3.2.
Figure 2(b) compares the strawman approach and TraCS-D. Assume the next location $\tau_{i+1}$ is a private location needing direction perturbation, thus the private direction is $\varphi:\tau_{i}^{\prime}\to\tau_{i+1}$. The dominant sector of TraCS-D is much smaller as $\varepsilon$ increases, e.g. $\varepsilon=6$ in Figure 2(b).  Specifically, when $\varepsilon=6$ and $\varphi=\pi/6$, it can be calculated that $p_{\varepsilon}\approx 3.20$ and $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})\approx[0.119\pi,0.214\pi)$. Thus, the dominant sector has a size of $0.095\pi\approx 17^{\circ}$ with probability $0.095\pi\times 3.20\approx 0.95$ of being chosen. In comparison, the strawman approach has a dominant sector of $2\pi/6=60^{\circ}$ with probability $0.98$ of being chosen.
Theoretical comparison between the strawman approach and Definition 3.1 is provided in Appendix B.2. This Appendix will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value.
3.2.2. Distance Perturbation
The other subspace of TraCS-D is the distance space $\mathcal{D}_{r(\varphi)}$. Besides the boundedness, the size of this space also varies with the reference location $\tau_{i}$ and the direction $\varphi$. The technical challenge in designing an LDP mechanism for this space lies in calculating the size $|\mathcal{D}_{r(\varphi)}|$ for any $\varphi$ at any $\tau_{i}$.
Given location $\tau_{i}$, any other location $\tau_{j}\in\mathcal{S}$ can be represented by a pair of coordinates $(\varphi,r(\varphi))$ relative to $\tau_{i}$. Since we focus on $\mathcal{S}$ as a rectangular area $[a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ has a closed-form expression depending on the direction $\varphi$. Specifically, this distance falls into one of the following four cases:
(1) $|\mathcal{D}_{r(\varphi)}|\in\left\{\frac{a_{\text{end}}-a_{i}}{\cos(\varphi)},\frac{b_{\text{end}}-b_{i}}{\sin(\varphi)},\frac{a_{i}-a_{\text{sta}}}{-\cos(\varphi)},\frac{b_{i}-b_{\text{sta}}}{-\sin(\varphi)}\right\},$
depending on the value of $\varphi$. Appendix B.3 provides the detailed equation form of Equation (1). Figure 3(b) illustrates the first two cases of $|\mathcal{D}_{r(\varphi)}|$. When $\varphi$ falls into $[0,\varphi_{1})$, i.e. Figure 3(a), $|\mathcal{D}_{r(\varphi)}|=(a_{\text{end}}-a_{i})/\cos(\varphi)$. Figure 3(b) shows the second case, i.e. when $\varphi\in[\varphi_{1},\varphi_{2})$, $|\mathcal{D}_{r(\varphi)}|=(b_{\text{end}}-b_{i})/\sin(\varphi)$.
(a) $\varphi\in[0,\varphi_{1})$
(b) $\varphi\in[\varphi_{1},\varphi_{2})$
Figure 3. Two cases of $|\mathcal{D}_{r(\varphi)}|$ at $\tau_{i}$ (blue lines).
[…]
Two decomposition of location space $S$ at $\tau_{i}$.
With the known distance space $[0,|\mathcal{D}_{r(\varphi)}|]$ and private distance $r(\varphi)$, we can design a distance perturbation mechanism to guarantee LDP for $\mathcal{D}_{r(\varphi)}$. For the convenience of presentation, we normalize the distance $r(\varphi)$ by $|\mathcal{D}_{r(\varphi)}|$, resulting in a normalized distance $\overline{r}(\varphi)\in[0,1)$.111 We use $[0,1)$ instead of $[0,1]$ to ease the presentation of Definition 3.3 and to align with the circular domain. These two domains are equivalent in implementation. Then we design a piecewise-based mechanism for the normalized distance $\overline{r}(\varphi)$.
Definition 3.3 (Distance Perturbation Mechanism).
Given a private distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, TraCS-D’s distance perturbation mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ is defined by:
$pdf[\mathcal{M}_{-}(\overline{r}(\varphi))=\overline{r}^{\prime}(\varphi)]=\begin{dcases}p_{\varepsilon}&\text{if}\ \overline{r}^{\prime}(\varphi)\in[u,v),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=\exp(\varepsilon/2)$ is the sampling probability, and $[u,v)$ is the sampling interval that
$\begin{split}[u,v)&=\begin{cases}\overline{r}(\varphi)+[-C,C)&\text{if}\ \overline{r}(\varphi)\in[C,1-C),\cr[0,2C)&\text{if}\ \overline{r}(\varphi)\in[0,C),\cr[1-2C,1)&\text{otherwise},\end{cases}\end{split}$
with $C=(\exp(\varepsilon/2)-1)/(2\exp(\varepsilon)-2)$.
Similar to $\mathcal{M}_{\circ}$, the above mechanism $\mathcal{M}_{-}$ has a higher probability density $p_{\varepsilon}$ in the central piece $[u,v)$, and a larger $\varepsilon$ results in a higher $p_{\varepsilon}$ and a narrower $[u,v)$, boosting the utility. It ensures LDP for $[0,1)$ and outputs a perturbed normalized distance $\overline{r}^{\prime}(\varphi)$. To cooperate with the direction perturbation, $\overline{r}^{\prime}(\varphi)$ can be mapped to another distance space $\mathcal{D}_{r(\varphi^{\prime})}$ at the perturbed direction $\varphi^{\prime}$. This is a linear mapping without randomness, so the post-processing property (Dwork and Roth, 2014) ensures it preserves the same privacy level.
Input: Location space $\mathcal{S}$, private trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$
Output: Perturbed trajectory $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$
$\mathcal{T}^{\prime}\leftarrow\emptyset,\mathcal{T}\leftarrow\tau_{0}^{\prime}\cup\mathcal{T}$;
[…]
1 for $i\leftarrow 0$ to $n-1$ do
2       $\tau_{i}=(a_{i},b_{i}),\tau_{i+1}=(a_{i+1},b_{i+1})$;
       $\varphi\leftarrow\text{atan2}(b_{i+1}-b_{i},a_{i+1}-a_{i})$;
[…]
       $\varphi^{\prime}\leftarrow\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$;
[…]
3       $R\leftarrow|\mathcal{D}_{r(\varphi)}|$ in Equation (1);
       $\overline{r}(\varphi)\leftarrow\frac{a_{i+1}-a_{i}}{R\cos(\varphi)}$;
[…]
       $\overline{r}^{\prime}(\varphi)\leftarrow\mathcal{M}_{-}(\overline{r}(\varphi);\varepsilon-\varepsilon_{d})$;
[…]
       $R^{\prime}\leftarrow|\mathcal{D}_{r(\varphi^{\prime})}|$, $r^{\prime}(\varphi^{\prime})\leftarrow\overline{r}^{\prime}(\varphi)\times R^{\prime}$;
[…]
4       $\tau_{i+1}^{\prime}=(a_{i}+r^{\prime}(\varphi^{\prime})\cos(\varphi^{\prime}),b_{i}+r^{\prime}(\varphi^{\prime})\sin(\varphi^{\prime}))$;
5       $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
       $\tau_{i+1}\leftarrow\tau_{i+1}^{\prime}$;
[…]
6
7return $\mathcal{T}^{\prime}$;
Algorithm 1 TraCS-D
3.2.3. Workflow of TraCS-D
Combining the direction perturbation and distance perturbation, TraCS-D perturbs each private location in the trajectory.
Algorithm 1 presents the workflow of TraCS-D. It takes the location space $\mathcal{S}$, the private trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input. Since TraCS-D relies on a non-private reference location, we add a dummy location $\tau_{0}^{\prime}$ to $\mathcal{T}$, which can be the starting coordinate of $\mathcal{S}$ or a randomly drawn location. Then, TraCS-D iterates over each pair of consecutive locations $\tau_{i}$ and $\tau_{i+1}$. For each private location $\tau_{i+1}$, it perturbs the direction (red block), and normalized distance (blue block) with privacy parameters $\varepsilon_{d}$ and $\varepsilon-\varepsilon_{d}$, respectively. The perturbed distance is then mapped along the perturbed direction (line 9-10). Line 12 updates the reference location to the perturbed location $\tau_{i+1}^{\prime}$ and performs the next iteration. Finally, it outputs the perturbed trajectory $\mathcal{T}^{\prime}$.
Algorithm 1 uses each $\tau_{i}^{\prime}$ as the reference location for the next perturbation. In fact, the reference location can be any non-private location in $\mathcal{S}$. We choose different $\tau_{i}^{\prime}$ for the consideration of alleviating the impact of a specific reference location.
3.2.4. Analysis of TraCS-D
This subsection analyzes the privacy, computational complexity, and utility of TraCS-D.
Theorem 3.4.
TraCS-D (Algorithm 1) satisfies $n\varepsilon$-LDP for the rectangular location space $\mathcal{S}$.
Proof.
(Sketch) The direction perturbation mechanism $\mathcal{M}_{\circ}$ satisfies $\varepsilon_{d}$-LDP by its definition. For the distance perturbation, the randomness comes entirely from mechanism $\mathcal{M}_{-}$, which satisfies $(\varepsilon-\varepsilon_{d})$-LDP. The post-processing property preserves the same privacy level after linearly mapping the perturbed distance. Then each location satisfies $\varepsilon$-LDP due to Composition Theorem 2.2. Hence, the entire trajectory perturbed by TraCS-D satisfies $n\varepsilon$-LDP. Appendix A.1 provides the detailed proof. ∎
Complexity. TraCS-D (Algorithm 1) has $\Theta(n)$ time complexity and $\Theta(n)$ space complexity, where $n$ is the length of the trajectory.  The time complexity is $\Theta(n)$ because each line of Algorithm 1 is $\Theta(1)$. The space complexity is $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$.
The $\Theta(n)$ time complexity is optimal for per-location perturbation mechanisms. Among existing LDP mechanisms for discrete location spaces, even if ignore the complexity of the score function and sampling in the Exponential mechanism (Nissim et al., 2012), NGram (Cunningham et al., 2021b) has $\Theta(tn)$ time complexity, where $t$ is the number of predefined “gram”. It also includes a large constant factor due to its search to satisfy reachability constraints. ATP (Zhang et al., 2023) is $\Theta(mn)$, where $m$ is the size of the location space, due to its necessary trajectory merging step. LDPTrace (Du et al., 2023) has $\Theta(mn)$ time complexity, where $m$ is the number of grids; it also has a large constant factor due to the model training procedure on a set of trajectories.
Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can be reduced to $\Theta(1)$ by updating each perturbed location $\tau_{i+1}^{\prime}$ in place within $\mathcal{T}$. This results in a lightweight memory footprint, which is crucial for mobile devices and edge computing.
Theorem 3.5.
In TraCS-D, both the worst-case mean square error (MSE) of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ converge to zero with a rate of $\Theta(e^{-\varepsilon/2})$.
Proof.
(Sketch) The MSE can be calculated by the closed-form expression of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$, allowing us to prove the convergence rate. Appendix A.2 provides the detailed proof. ∎
This theorem indicates that both the error of the perturbed direction and distance are reduced exponentially with the privacy parameter $\varepsilon$.  Note that the size of the distance space also affects the variance. Denote the output of $\mathcal{M}_{-}$ as a random variable $Y$ and the private input as $x$. After the linear mapping, the perturbed distance is $Y\cdot|\mathcal{D}_{r}|$ and the private distance is $x\cdot|\mathcal{D}_{r}|$, where $|\mathcal{D}_{r}|$ is the size of the distance space. Thus, the MSE of the perturbed distance is $\text{MSE}[Y\cdot|\mathcal{D}_{r}|]=|\mathcal{D}_{r}|^{2}\cdot\text{MSE}[Y]$. Therefore, a smaller $|\mathcal{D}_{r}|$ leads to a smaller MSE of the perturbed distance.
3.3. TraCS-C
TraCS-C decomposes the rectangular location space $\mathcal{S}$ into two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, as shown in Figure 1(b).  Specifically, we can fix $(a_{\text{sta}},b_{\text{sta}})$ as the reference location, with longitude $a_{\text{sta}}$ as the $a$-axis and latitude $b_{\text{sta}}$ as the $b$-axis. Then any location $\tau_{i}\in\mathcal{S}$ can be represented as a pair of Cartesian coordinates $(d_{a},d_{b})$, where $d_{a}\in\mathcal{D}_{a}$ and $d_{b}\in\mathcal{D}_{b}$ are the distances from $\tau_{i}$ to the $a$-axis and $b$-axis, respectively. This representation is unique and reversible: any pair of distances $(d_{a},d_{b})$ can be mapped back to a unique location $\tau_{i}\in\mathcal{S}$.  Thus, if $\tau_{i}$ is a private location, an LDP mechanism applied to $\tau_{i}$ is essentially applied to $d_{a}$ and $d_{b}$. By Composition Theorem 2.2, we can achieve LDP for $\tau_{j}$ by applying perturbation mechanisms for $d_{a}$ and $d_{b}$, respectively.
3.3.1. Coordinates Perturbation
In TraCS-C’s representation of locations, $\mathcal{D}_{a}=[0,a_{\text{end}}-a_{\text{sta}})$ and $\mathcal{D}_{b}=[0,b_{\text{end}}-b_{\text{sta}})$ are independent of any specific location. Therefore, the location perturbation of any $\tau_{i}$ can be performed independently in $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$.
Following the same idea of distance perturbation in TraCS-D, we normalize the distance $d_{a}$ and $d_{b}$ to $[0,1)$ by dividing $|\mathcal{D}_{a}|$ and $|\mathcal{D}_{b}|$, respectively. Then the mechanism $\mathcal{M}_{-}$ in Definition 3.3 provides $\varepsilon$-LDP for the normalized distances, and the linear mappings back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ preserve the privacy guarantee.
Input: Location space $\mathcal{S}$, private trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$
Output: Perturbed trajectory $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$
1 $\mathcal{T}^{\prime}\leftarrow\emptyset$;
2 for $i\leftarrow 1$ to $n$ do
3       $\tau_{i}=(a_{i},b_{i})$;
       $(\overline{d}_{a},\overline{d}_{b})\leftarrow(\frac{a_{i}-a_{\text{sta}}}{a_{\text{end}}-a_{\text{sta}}},\frac{b_{i}-b_{\text{sta}}}{b_{\text{end}}-b_{\text{sta}}})$;
[…]
4       $\overline{d}^{\prime}_{a}\leftarrow\mathcal{M}_{-}(\overline{d}_{a};\varepsilon/2)$;
       $\overline{d}^{\prime}_{b}\leftarrow\mathcal{M}_{-}(\overline{d}_{b};\varepsilon/2)$;
[…]
       $(d_{a}^{\prime},d_{b}^{\prime})\leftarrow(\overline{d}^{\prime}_{a}|\mathcal{D}_{a}|,\overline{d}^{\prime}_{b}|\mathcal{D}_{b}|)$;
[…]
5       $\tau^{\prime}_{i}=(a_{\text{sta}},b_{\text{sta}})+(d_{a}^{\prime},d_{b}^{\prime})$;
6       $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{\tau^{\prime}_{i}\}$;
7
8return $\mathcal{T}^{\prime}$;
Algorithm 2 TraCS-C
Algorithm 2 shows the workflow of TraCS-C. It takes the location space $\mathcal{S}$, the private trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input, and outputs the perturbed trajectory $\mathcal{T}^{\prime}$. For each location $\tau_{i}=(a_{i},b_{i})$ in $\mathcal{T}$, Line 4 computes its coordinates $(d_{a},d_{b})$ and normalizes them to $(\overline{d}_{a},\overline{d}_{b})\in[0,1)\times[0,1)$. Lines 5-6 perturb the normalized coordinates $\overline{d}_{a}$ and $\overline{d}_{b}$ using $\mathcal{M}_{-}$ with privacy parameter $\varepsilon/2$. Then Line 7 maps the perturbed normalized coordinates back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, and Line 8 converts them to the longitude-latitude representation $\tau^{\prime}_{i}$.
3.3.2. Analysis of TraCS-C
Theorem 3.6.
TraCS-C (Algorithm 2) satisfies $n\varepsilon$-LDP for the rectangular location space $\mathcal{S}$.
Proof.
(Sketch) Algorithm 2 uses mechanism $\mathcal{M}_{-}$ twice, each with a privacy parameter $\varepsilon/2$. By Composition Theorem 2.2, their composition satisfies $\varepsilon$-LDP. The subsequent linear mappings to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ are post-processing steps that preserve the same privacy. Hence, each perturbed location satisfies $\varepsilon$-LDP, and the whole trajectory satisfies $n\varepsilon$-LDP. Appendix A.3 provides the details. ∎
Complexity. TraCS-C (Algorithm 2) has $\Theta(n)$ time complexity and $\Theta(n)$ space complexity, where $n$ is the length of the trajectory.  The time complexity is $\Theta(n)$ because each line of Algorithm 2 is $\Theta(1)$. The space complexity is $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$. Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can also be reduced to $\Theta(1)$ by updating each perturbed location $\tau_{i}^{\prime}$ in place within $\mathcal{T}$.
3.4. Rounding to Discrete Space
Although TraCS is designed for continuous location spaces, it can be applied to any fine-grained discrete location space by rounding the perturbed locations to the nearest discrete locations. The rounding is a post-processing step that does not affect the privacy guarantee for the continuous space.
Specifically, if the discrete location space is a set of cells, we can round the perturbed location to the cell that contains it. Formally, assume $\mathcal{C}=\{c_{1},\ldots,c_{m}\}\subseteq\mathcal{S}$ discretize $\mathcal{S}$ into $m$ cells. If $\tau_{i}^{\prime}$ is the perturbed location of $\tau_{i}$ in TraCS, then the discretized TraCS outputs $c_{j}$ such that $\tau_{i}^{\prime}\in c_{j}$. Commonly, the cells are evenly divided, so the rounding costs $\Theta(1)$ time, as $c_{j}$ can be determined by the coordinates of $\tau_{i}^{\prime}$.
If the location space is a set of location points, we can round the perturbed location to the nearest location point. Formally, assume $\mathcal{P}=\{p_{1},\ldots,p_{m}\}\subseteq\mathcal{S}$ is a set of location points, and $\tau_{i}^{\prime}$ is the perturbed location of $\tau_{i}$ in TraCS. Then, the discretized TraCS outputs $\operatorname*{arg\,min}_{p_{j}\in\mathcal{P}}||\tau_{i}^{\prime}-p_{j}||_{2}$. The time complexity of rounding to the nearest location point is $\mathcal{O}(m)$ in the worst case. In practice, this time complexity can be significantly reduced by using heuristic or approximation algorithms.
3.5. Discussion and Extensions
This subsection discusses the comparison between TraCS-D and TraCS-C, and their extensions to other types of location spaces.
3.5.1. Privacy Comparison Between TraCS-D and TraCS-C
Although TraCS-D and TraCS-C provide the same level of privacy guarantee, they have different privacy interpretations.
In TraCS-D, the LDP guarantee is provided for the direction information and the subsequent distance information. This means that when the usage of TraCS-D and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the private direction $\varphi:\tau^{\prime}_{i}\to\tau_{i+1}$ and the private distance $r(\varphi):|\tau^{\prime}_{i}\to\tau_{i+1}|$ from the known $\tau^{\prime}_{i}$ and $\tau^{\prime}_{i+1}$.  Although the distance space $\mathcal{D}_{r(\varphi)}$ relies on a specific direction, it does not leak the direction information, as the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of any location and TraCS-D uses the perturbed direction, which preserves LDP.
In TraCS-C, the LDP guarantee is provided for the two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$. This means that when the usage of TraCS-C and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the private distances $d_{a}$ and $d_{b}$.
3.5.2. Impact of the Size of $\mathcal{S}$
Intuitively, a larger location space $\mathcal{S}$ requires larger privacy parameters to achieve the same level of utility. For TraCS-D and TraCS-C, this impact is reflected in the size of the direction space and distance spaces.
In TraCS-D, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of the size of $\mathcal{S}$, meaning that the MSE of the perturbed direction, $\mathrm{MSE}[\varphi^{\prime}]$, remains constant regardless of the size of $\mathcal{S}$.  The distance space $\mathcal{D}_{r(\varphi)}$ depends on the specific direction $\varphi$ and the size of $\mathcal{S}$. In the worst case, $\mathcal{D}_{r(\varphi)}$ is the diagonal of $\mathcal{S}$. We have shown that the MSE of $\mathcal{M}_{-}$ is quadratic to the size of the distance space after linear mapping in Section 3.2.4. Therefore, when $|\mathcal{D}_{r(\varphi)}|$ increases linearly, the expected error of the perturbed distance, i.e. $\mathrm{E}[|r^{\prime}(\varphi)-r(\varphi)|]$, also increases linearly.  However, the error in the direction space $\mathcal{D}_{\varphi}$ affects the error in $\mathcal{D}_{r(\varphi)}$ in a non-linear way. Specifically, under the Euclidean distance metric, this effect is $|\mathcal{D}_{r(\varphi)}|\cdot 2\sin(|\varphi-\varphi^{\prime}|/2)$, which is the chord length of the arc between $\varphi$ and $\varphi^{\prime}$.
In TraCS-C, both $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ are determined by the size of $\mathcal{S}$. Since they use mechanism $\mathcal{M}_{-}$ independently, their expected error increases linearly with the increase of $|\mathcal{D}_{a}|$ and $|\mathcal{D}_{b}|$. Meanwhile, the error in $\mathcal{D}_{a}$ affects the error in $\mathcal{D}_{b}$ linearly.
3.5.3. Other Piecewise-based Mechanisms
In Definition 3.1 and Definition 3.3, we use specific instantiations for the parameters $p$ and $[l,r)$ in the piecewise-based mechanism. Besides these instantiations, TraCS can be integrated with any other piecewise-based mechanisms (Wang et al., 2019; Li et al., 2020; Ma et al., 2024; Liu et al., 2024; Zhao et al., 2021) by redesigning them. For example, SW (Li et al., 2020) is a famous mechanism proposed for distribution estimation; it is defined on $[0,1)\to[-b,1+b)$, with $b$ as a parameter. It uses different $p$ and $[l,r)$, and employs a strategy of maximizing the mutual information between the perturbed and private data to determine $b$.  We can redesign SW for the direction perturbation and distance perturbation in TraCS-D and TraCS-C. Appendix B.4 provides the detailed redesign of SW for them and shows intuitive comparisons with Definition 3.1 and Definition 3.3.
3.5.4. Other Shapes of Location Space
The design of LDP mechanisms is space-specific, implying that extensions to other shapes of location space are non-trivial.
For TraCS, the challenge lies in determining the size of each subspace, i.e. $|\mathcal{D}_{\varphi}|$, $|\mathcal{D}_{r(\varphi)}|$, $|\mathcal{D}_{a}|$, and $|\mathcal{D}_{b}|$. Among them, $|\mathcal{D}_{r(\varphi)}|$ is not apparent but can be calculated for rectangular location spaces.  Although calculating $|\mathcal{D}_{r(\varphi)}|$ is generally non-trivial for irregular shapes, TraCS can be extended to parallelogram location spaces by a linear transformation to a rectangular space. Meanwhile, TraCS-D can be used for circular areas, which are widely used in relative location representation, e.g. radar and sonar systems.
TraCS-D for circular area. When the location space is circular with radius $R$, TraCS-D has a more straightforward design. The main insight is that such a location space $\mathcal{S}_{R}$ has a natural representation in polar coordinates, i.e. $\mathcal{S}_{R}=\{(\varphi,r):\varphi\in[0,2\pi),r\in[0,R)\}$. Any location $\tau_{i}\in\mathcal{S}_{R}$ corresponds to a pair of coordinates $(\varphi_{i},r_{i})$, and this representation is unique and reversible.  In this case, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ and the distance space $\mathcal{D}_{r}=[0,R)$ are independent, so we can apply $\mathcal{M}_{\circ}$ to $\varphi_{i}$ and $\mathcal{M}_{-}$ to $r_{i}/R$.
4. Evaluation
This section evaluates the performance of TraCS 222 Code, experiments, and slides are available at https://github.com/ZhengYeah/TraCS. on both continuous and discrete spaces.
4.1. Results on Continuous Space
The strawman method, designed for direction perturbation, can be combined with the same distance perturbation in TraCS-D to ensure LDP for continuous space $\mathcal{S}\subset\mathbb{R}^{2}$. We will still refer to this extended method as the strawman method and compare it with TraCS in continuous space. Additionally, we will examine the effects of different piecewise-based mechanisms and evaluate the performance of TraCS in circular areas.
4.1.1. Setup
(a) $\mathcal{S}=[0,1)\times[0,1)$
(b) $\mathcal{S}=[0,2)\times[0,10)$
Figure 4. Comparison of TraCS and the strawman method.
[…]
Comparison of TraCS and the strawman method.
Figure 5. Conditions TraCS-D outperforms TraCS-C.
Figure 6. TraCS-D vs $k$-RR + uniform sampling.
[…]
TraCS-D vs $k$-RR + uniform sampling.
TraCS requires the following parameters: the location space $\mathcal{S}$, the private trajectory $\mathcal{T}$, the privacy parameter $\varepsilon$, and $\varepsilon_{d}$ for direction perturbation in TraCS-D. The strawman method also requires an instance of $k$ in $k$-RR. We use configurations:
  • Location space $\mathcal{S}$: The default location space is $[0,1)^{2}$. We consider a larger space $\mathcal{S}=[0,2)\times[0,10)$ for comparison.
  • Privacy parameter $\varepsilon_{d}$: We set $\varepsilon_{d}=\varepsilon\pi/(\pi+1)$ to heuristically balance the perturbation in $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ according to their domain sizes.
  • $k$-RR: We set $k=6$ as the strawman method’s default value, and consider $3$-RR and $12$-RR for comparison.
Trajectory utility metric. We follow the commonly used Euclidean distance to measure the utility of the perturbed trajectory. The distance between two locations $\tau_{i}$ and $\tau^{\prime}_{i}$ is defined as $\|\tau_{i}-\tau^{\prime}_{i}\|_{2}$. Therefore, given a private trajectory $\mathcal{T}$ and a perturbed trajectory $\mathcal{T}^{\prime}$, the average error (AE) among all locations in the trajectory is:
$AE(\mathcal{T},\mathcal{T}^{\prime})=\frac{1}{|\mathcal{T}|}\sum_{i=1}^{|\mathcal{T}|}\|\tau_{i}-\tau^{\prime}_{i}\|_{2},$
where $|\mathcal{T}|$ is the number of locations in the trajectory. A smaller $AE$ indicates better utility of the perturbed trajectory.
To eliminate the impact of specific trajectories, we generate $100$ random trajectories for each experiment and report the average $AE$ over all the trajectories.
4.1.2. Error Comparison
Figure 4(b) presents the error comparison between TraCS and the strawman method. We can observe the error advantage of TraCS over the strawman method as the privacy parameter $\varepsilon$ increases. Specifically, in Figure 4(a), TraCS-C exhibits smaller $AE$ when $\varepsilon\gtrsim 3$, and TraCS-D exhibits smaller $AE$ when $\varepsilon\gtrsim 5$. When the location space expands to $[0,2)\times[0,10)$, i.e. Figure 4(b), the advantage of TraCS becomes more significant, outperforming the strawman method at smaller $\varepsilon$ values.  Among TraCS, TraCS-C consistently has a smaller $AE$ than TraCS-D. This is because TraCS-D often deals with larger distance spaces $\mathcal{D}_{r(\varphi)}$, which also magnifies the error of direction perturbation.  Statistically, the mean $AE$ of TraCS-D is $91.1\%$ of the strawman method in Figure 4(a) and $86.6\%$ in Figure 4(b) across all $\varepsilon$ values. The mean $AE$ of TraCS-C is $75.5\%$ of the strawman method in Figure 4(a) and $64.0\%$ in Figure 4(b) across all $\varepsilon$ values.
Conditions TraCS-D outperforms TraCS-C. The error of TraCS-C depends on the size of the distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, while the error of TraCS-D depends only on the size of the distance space $\mathcal{D}_{r(\varphi)}$. Therefore, if $\mathcal{D}_{r(\varphi)}$ is majorly smaller than $\mathcal{D}_{a}$ or $\mathcal{D}_{b}$, TraCS-D will outperform TraCS-C.  We conduct experiments to verify this condition. Specifically, we set $\mathcal{S}=[0,2)\times[0,10)$, and choose a private trajectory along the $x$-axis, i.e. the short side of $\mathcal{S}$. Then we perturb the trajectory $1000$ times using TraCS-C and TraCS-D and calculate the average $AE$. In this case, $|\mathcal{D}_{r(\varphi)}|\approx 2$, much smaller than $|\mathcal{D}_{b}|=10$. Consequently, TraCS-D exhibits smaller $AE$ than TraCS-C across all the $\varepsilon$ values, as shown in Figure 6. In this experiment, the mean $AE$ of TraCS-D is $49.8\%$ of TraCS-C across all the $\varepsilon$ values.
Figure 7. Comparison between different piecewise-based mechanisms.
Figure 8. Example: TraCS-D for circular area.  Private location $\tau=(\pi,0.5)$.
Error of direction perturbation. Figure 6 shows the average error of direction perturbation in TraCS-D compared with the strawman method. We collect the private directions of the trajectories in the location space $\mathcal{S}=[0,1)\times[0,1)$, and perform direction perturbation using TraCS-D and the strawman method. For the strawman method, we use $k$-RR with $k=3,6,12$. This experiment reflects empirical error results of mechanism $\mathcal{M}_{\circ}$ in Definition 3.1 and $k$-RR + uniform sampling.  We can observe an exponentially decreasing error of TraCS-D as $\varepsilon$ increases, as stated in Theorem 3.5. In contrast, the strawman method exhibits an eventually stable error as $\varepsilon$ increases, due to the inherent inner-sector error of $k$-RR.
4.1.3. Effect of Different Piecewise-based Mechanisms
Figure 8 compares the performance of different piecewise-based mechanisms for TraCS-D. In addition to the mechanisms defined in Definition 3.1 and Definition 3.3, we name the redesigned SW for TraCS-D as TraCS-D (R-SW), as detailed in Appendix B.4.  We can observe that TraCS-D consistently exhibit smaller errors than TraCS-D (R-SW) across all the $\varepsilon$ values. Statistically, the mean $AE$ of TraCS-D is $52.6\%$ of TraCS-D (R-SW) across all the $\varepsilon$ values. This difference comes from the MSE of their sampling distributions, where Definition 3.1 and Definition 3.3 have smaller MSE than R-SW.
4.1.4. TraCS-D for Circular Area
Figure 8 illustrates an example of TraCS-D applied to a circular area $\mathcal{S}=[0,2\pi)\otimes[0,1)$. We set the private location $\tau$ as $(\pi,0.5)$ and the privacy parameter $\varepsilon=5$, then collect $50$ random samples of perturbed locations using TraCS-D. In this case, the dominant area is $[0.87\pi,1.13\pi)\otimes[0.32,0.67)$, as determined by the perturbation mechanisms in Definition 3.1 and Definition 3.3. We can observe that the perturbed locations are concentrated in the dominant area.
4.1.5. Choose TraCS-D or TraCS-C
From the above experiments, we can conclude criteria for choosing TraCS-D or TraCS-C: (i) TraCS-C generally has better trajectory utility than TraCS-D for random trajectories; (ii) TraCS-D is better for specific trajectories where their distance spaces are smaller than in TraCS-C, and suitable for circular areas.
4.2. Results on Discrete Space
TraCS can be applied to discrete spaces by rounding the perturbed locations to the nearest discrete points. Notably, the privacy guarantee remains valid for the continuous space and hence is also valid for the discrete space, independent of the discretization. We will evaluate the performance of TraCS on discrete spaces and compare it with existing methods.
4.2.1. Setup
(a) $m=10\times 10$
(b) $m=50\times 50$
Figure 9. Comparison on synthetic datasets. $m$ is the number of discrete points in the location space.
[…]
Comparison of TraCS and the strawman method.
We compare the following methods:
  • ATP (Zhang et al., 2023): We set the $k$-RR parameter $k=6$ in ATP. Although it uses a trajectory region constraining strategy to reduce the location space for perturbation, we omit this strategy and instead use reachability constraint in NGram to illustrate the effect.
  • NGram (Cunningham et al., 2021b): We consider a strong reachability constraint, i.e., from each location, the next location can only be within a distance of $0.75\cdot\hat{a}$, where $\hat{a}$ is the distance between the start and end longitude of the location space. This constraint significantly reduces the location space for perturbation, especially near the boundary.
  • TraCS: We use the same configurations as in the continuous space experiments. For each perturbed location from TraCS, we round it to the nearest discrete location.
The privacy parameter $\varepsilon$ in ATP is for an entire trajectory and can be distributed to each location. For a fair comparison, we use $\varepsilon$ as the privacy parameter for each location. NGram with reachability constraint also uses the same $\varepsilon$ for each location. We omit LDPTrace (Du et al., 2023) as it focuses on trajectory synthesis for a location space rather than perturbation for a given trajectory. The comparison is conducted on the following datasets:
  • Synthetic dataset: We discretize $[0,1)\times[0,1)$ into $m$ discrete points and treat the $m$ points as the location space.
  • NYC and CHI: These real-world datasets are the same as in ATP, extracted from the Foursquare dataset (Yang et al., 2015) and Gowalla dataset (Cho et al., 2011), respectively. [^3] NYC is in $[-74.26,-73.70]\times[40.56,40.97]$ and CHI in $[-87.9,-87.5]\times[41.6,42.0]$. NYC contains $3,478$ discrete locations; CHI contains $1,000$ discrete locations. We use the first $100$ trajectories for evaluation.
For the synthetic dataset, we generate $100$ random trajectories with $100$ locations each to compute the average $AE$. For the NYC and CHI datasets, we perform the perturbation $5$ times for each trajectory and compute the average $AE$.
(a) NYC dataset
(b) CHI dataset
Figure 10. Comparison on real-world datasets.
[…]
Comparison on real-world datasets.
Table 2. Time cost comparison.
ATP NGram TraCS-D TraCS-C
Total (ms) $145.7$ $100.9$ $0.06$ $0.05$
Perturbation $125.8$ $92.8$ $0.018$ $0.003$
4.2.2. Error Comparison on Synthetic Dataset
Figure 9(b) presents the error comparison on the synthetic dataset. Figure 9(a) shows results for $m=10\times 10$, and Figure 9(b) shows results for $m=50\times 50$. We observe that TraCS-C consistently outperforms ATP and NGram across all the $\varepsilon$ values, and TraCS-D also outperforms ATP and NGram when $\varepsilon\gtrsim 4$. Moreover, the performance advantage of TraCS becomes more significant as $m$ increases. This is because finer-grained discretization allows TraCS’s rounding to align with the private locations more accurately. Statistically, the mean $AE$ of TraCS-C is $65.2\%$ of ATP and $77.4\%$ of NGram in Figure 9(a), and $61.5\%$ of ATP and $73.2\%$ of NGram in Figure 9(b). The mean $AE$ of TraCS-D is $75.5\%$ of ATP and $89.6\%$ of NGram in Figure 9(a), and $89.0\%$ of ATP and $74.7\%$ of NGram in Figure 9(b). These results indicate the advantage of TraCS for synthetic trajectories in discrete spaces.
4.2.3. Error Comparison on Real-world Dataset
Figure 10(b) presents the error comparison on the NYC and CHI datasets. ATP and NGram perform better than in the synthetic dataset, especially when $\varepsilon$ is small. However, TraCS still outperforms them as $\varepsilon$ increases.
Reasons for the results. (i) ATP performs exceptionally well when $\varepsilon$ is small, particularly on the NYC dataset. This is because this dataset contains many “condensed” locations, i.e., no other locations near the trajectory. In such cases, if the bi-direction perturbation of ATP is accurate, the perturbed location is error-free. (ii) The $AE$ of ATP does not decrease significantly as $\varepsilon$ increases, which is consistent with their paper. NGram shows a similar trend, especially on the NYC dataset. This is due to the low performance of the Exponential mechanism for large location spaces, particularly when the locations are concentrated in a small area. In such cases, the score functions of the mechanism struggle to differentiate the perturbation quality among the locations. (iii) TraCS exhibits larger $AE$ than ATP and NGram when $\varepsilon$ is small because it focuses on the rectangular area formed by the longitude and latitude, instead of a set of discrete locations. Therefore, the perturbed locations may be far from the city center when $\varepsilon$ is small.
4.2.4. Time Cost Comparison
Table 2 presents the time cost comparison on the NYC dataset. The time cost is measured in milliseconds and averaged per location. It shows that TraCS-D and TraCS-C are significantly faster than ATP and NGram. This efficiency is primarily due to the piecewise-based mechanisms in TraCS, which require negligible time for perturbation compared to the Exponential mechanism in ATP and NGram.
ATP has the highest time cost because it processes two copied trajectories. NGram has a lower time cost than ATP but is still significantly higher than TraCS. These high time costs are primarily due to the Exponential mechanism, which requires multiple evaluations of score functions and sampling from complex distributions (detailed in Appendix B.1). The perturbation procedure accounts for $86.3\%$ and $92.0\%$ of the total time cost in ATP and NGram, respectively.
The total time cost of TraCS is less than $0.5\%$ of ATP and NGram. The perturbation procedure of TraCS-D and TraCS-C only takes $30\%$ and $6\%$ time, respectively. TraCS-D has a higher time cost than TraCS-C because it needs to calculate the distance space $\mathcal{D}_{r(\varphi)}$ for each location in the trajectory, while TraCS-C can use the same $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ for all locations in the trajectory.
5. Related Work
Trajectory Collection under LDP
Based on the privacy notion, trajectory collection methods under LDP can be categorized into two groups: pure LDP and relaxed LDP. Compared with relaxed LDP, pure LDP imposes a stronger constraint for indistinguishability.
Existing trajectory collection methods under pure LDP include NGram (Cunningham et al., 2021b), LDPTrace (Du et al., 2023), and ATP (Zhang et al., 2023). Among them, LDPTrace focuses on trajectory synthesis within a given location space, while NGram, ATP, and this paper focus on trajectory perturbation.  NGram attempts to ensure pure LDP on the “trajectory space”, which is stronger than the location space. However, this is proved extremely challenging as the trajectory space increases exponentially with the number of possible locations. ATP and this paper focus on ensuring pure LDP in the location space. Unlike ATP and other existing methods, this paper focuses on continuous location spaces, which are stronger than any fine-grained discrete location space.
As a special case of trajectory collection, L-SRR (Wang et al., 2022) focuses on single-location publishing. It discretizes the location space into different-sized cells, with smaller cells near the private location and larger cells farther away. Then, SRR (a variant of $k$-RR) is applied to the cells. Similar to other discretization methods, the number of cells affects its privacy and utility. TraCS provides solutions to this problem in continuous spaces, achieving better privacy and utility, particularly when the number of cells is large.
A widely used relaxed LDP notion for trajectory collection is Geo-indistinguishability (Andrés et al., 2013; Mendes et al., 2020), where the indistinguishability between two locations is a function of their distance, i.e. privacy decreases with distance. Input-discriminative LDP (Gu et al., 2020) goes one step further, assigning each location its privacy level. Threshold-integrated LDP (Yang et al., 2024) defines a threshold to constrain the private region near a location. This is similar to NGram’s reachability constraint but with a formal privacy definition.  More relaxed LDP notions for trajectory collection can be found in (Miranda-Pascual et al., 2023). Although these relaxed LDP notions generally provide better trajectory utility than pure LDP, they sacrifice the strong privacy guarantee of pure LDP, and often need hyperparameters to define the privacy level.
Trajectory Collection under DP
As a weaker privacy notion than LDP, DP (Dwork, 2006; Dwork and Roth, 2014) requires only that any two neighboring locations in the location space be indistinguishable, rather than any two arbitrary locations.
The usage of direction information in trajectory collection dates back to an early work SDD (Jiang et al., 2013), which attempts to add unbounded Laplace noise to the direction information. It provides a DP guarantee for unbounded continuous location spaces, but results in poor utility. The notion of “n-gram” dates back to (Chen et al., 2012), where it abstracts a trajectory pattern as a gram.
External knowledge is also widely used in DP-based trajectory collection methods. NGram’s prior work (Cunningham et al., 2021a) provides a DP-based solution for trajectory synthesis by exploiting the public knowledge of the road network. Focusing on particular privacy-concerned locations, or sensitive zones (Wang and Kankanhalli, 2020), can also improve the trajectory utility while maintaining partial privacy guarantees.
6. Conclusions
This paper presents TraCS-D and TraCS-C, providing local differential privacy guarantees for trajectory collection in continuous location spaces. TraCS-D uses a novel direction-distance perturbation procedure, while TraCS-C perturbs the Cartesian coordinates of each location. These two methods can also be applied to discrete location spaces by rounding the perturbed location to the nearest discrete point. Their trajectory utility is theoretically analyzed and empirically evaluated. Evaluation results on discrete location spaces show that TraCS outperforms state-of-the-art methods in trajectory utility, particularly when the privacy parameter is large.
Beyond trajectory collection, the proposed perturbation mechanisms can also be applied to other scenarios with continuous and bounded domains, providing a direction for future work.
Appendix A Proofs
A.1. Proof of Theorem 3.4
Proof.
Denote TraCS-D in Algorithm 1 as $\mathcal{M}$. We need to prove that $\mathcal{M}$ satisfies $\varepsilon$-LDP for each location $\tau\in\mathcal{S}$, i.e.
$\forall\tau_{1},\tau_{2},\tau^{*}\in\mathcal{S}:\frac{pdf[\mathcal{M}(\tau_{1})=\tau^{*}]}{pdf[\mathcal{M}(\tau_{2})=\tau^{*}]}\leq\exp(\varepsilon).$
In TraCS-D, each location $\tau$ is represented in a direction-distance space with unique coordinates $(\varphi,r(\varphi))$. The key insight of the proof is that the perturbation of $\varphi$ and $r(\varphi)$ satisfies $\varepsilon_{d}$-LDP and $(\varepsilon-\varepsilon_{d})$-LDP, respectively. Then TraCS-D ensures $\varepsilon$-LDP for the location space $\mathcal{S}=\mathcal{D}_{\varphi}\otimes\mathcal{D}_{r(\varphi)}$.
The definitions of mechanisms $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ imply that they satisfy LDP. Specifically, for $\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$,
$\forall\varphi_{1},\varphi_{2},\varphi^{*}\in[0,2\pi):\frac{pdf[\mathcal{M}_{\circ}(\varphi_{1})=\varphi^{*}]}{pdf[\mathcal{M}_{\circ}(\varphi_{2})=\varphi^{*}]}\leq\exp(\varepsilon_{d}),$
is guaranteed by Definition 3.1 of $\mathcal{M}_{\circ}$. Similarly, $\mathcal{M}_{-}(\overline{r}(\varphi);\varepsilon-\varepsilon_{d})$ in definition 3.3 ensures $(\varepsilon-\varepsilon_{d})$-LDP for $\overline{r}(\varphi)$, i.e.
$\forall\overline{r}_{1},\overline{r}_{2},\overline{r}^{*}\in[0,1):\frac{pdf[\mathcal{M}_{-}(\overline{r}_{1})=\overline{r}^{*}]}{pdf[\mathcal{M}_{-}(\overline{r}_{2})=\overline{r}^{*}]}\leq\exp(\varepsilon-\varepsilon_{d}).$
The subsequent linear mapping from $\overline{r}\in[0,1)$ to $r\in\mathcal{D}_{r(\varphi^{\prime})}$ is a post-processing step that does not affect the LDP property. Specifically, denote the linear mapping as $g:[0,1)\to\mathcal{D}_{r(\varphi^{\prime})}$, and denote $r_{1}=g(\mathcal{M}_{-}(\overline{r}_{1}))$ and $r_{2}=g(\mathcal{M}_{-}(\overline{r}_{2}))$ as two random variables. It is clear that
$\forall r_{1},r_{2},r^{*}\in\mathcal{D}_{r(\varphi^{\prime})}:\frac{pdf[r_{1}=r^{*}]}{pdf[r_{2}=r^{*}]}\leq\exp(\varepsilon-\varepsilon_{d}),$
because the linear mapping $g$ is a deterministic function without randomness. Therefore, $\mathcal{M}_{-}$ ensures $(\varepsilon-\varepsilon_{d})$-LDP for any $\mathcal{D}_{r(\varphi^{\prime})}$.
Combining these results with Sequential Composition Theorem 2.2, we can conclude that $\mathcal{M}=(\mathcal{M}_{\circ},\mathcal{M}_{-})$ ensures $\varepsilon$-LDP for each location $\tau=(\varphi,r(\varphi))\in\mathcal{S}$. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
A.2. Proof of Theorem 3.5
Proof.
We prove the $\Theta(e^{-\varepsilon/2})$ convergence rate by calculating the MSE of the perturbation mechanism. Specifically, the MSE of $\mathcal{M}_{-}$ is calculated as follows:
[…]
The worst-case MSE is achieved when $\overline{r}=0$ or $\overline{r}=1$, i.e. at the endpoints (similar to the variance calculation in (Wang et al., 2019)). In this case,
[…]
with $p_{\varepsilon}$ and $C$ defined in Definition 3.3. We omit the calculation for $\overline{r}=1$ as it is symmetric to $\overline{r}=0$. Plugging in the values of $p_{\varepsilon}$ and $C$, we can simplify the MSE using big $\mathcal{O}$ notation:
$\begin{split}\mathrm{MSE}[\mathcal{M}_{-}(0)]&=\frac{1}{3}e^{-\varepsilon/2}+\frac{(e^{\varepsilon/2}-1)^{3}}{3e^{\varepsilon/2}(e^{\varepsilon}-1)^{2}}\\ &=\frac{1}{3}e^{-\varepsilon/2}+\mathcal{O}(e^{-\varepsilon}).\end{split}$
Therefore, as $\varepsilon\to\infty$, the worst-case MSE of $\mathcal{M}_{-}$ converges to zero at a rate of $\Theta(e^{-\varepsilon/2})$. The MSE of $\mathcal{M}_{\circ}$ can be calculated in the same way, resulting in a convergence rate of $\Theta(e^{-\varepsilon/2})$ as well. ∎
A.3. Proof of Theorem 3.6
Proof.
Denote TraCS-C in Algorithm 2 as $\mathcal{M}$. We need to prove that $\mathcal{M}$ satisfies $\varepsilon$-LDP for each location $\tau\in\mathcal{S}$. In TraCS-C, each location $\tau$ is represented in Cartesian space $\mathcal{D}_{a}\times\mathcal{D}_{b}$ with unique coordinates $(a,b)$. Therefore, the key insight is to prove that the perturbation of $a$ and $b$ satisfies $\varepsilon/2$-LDP each.
In Proof A.1, we have already proved that $\mathcal{M}_{-}(\overline{a},\varepsilon/2)$ satisfies $\varepsilon/2$-LDP for the normalized distance $\overline{a}$. Then, the linear mapping from $\overline{a}\in[0,1)$ to $a\in\mathcal{D}_{a}$ is a post-processing step that does not affect the LDP property. This means $\mathcal{M}_{-}$ ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{a}$. In the same way, it also ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{b}$. Therefore, $\mathcal{M}=(\mathcal{M}_{-},\mathcal{M}_{-})$ satisfies $\varepsilon$-LDP for each location $\tau=(a,b)\in\mathcal{S}$ by Sequential Composition Theorem 2.2. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
Appendix B Complementary Materials
B.1. Weakness of the Exponential Mechanism
This section discusses the Exponential mechanism and its weaknesses in the context of trajectory collection.
Definition B.1 (Exponential Mechanism (Nissim et al., 2012)).
Given a scoring function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, a privacy parameter $\varepsilon>0$, and a set of possible outputs $\mathcal{Y}$, the Exponential mechanism $\mathcal{M}_{\exp}:\mathcal{X}\to\mathcal{Y}$ is defined by:
$\Pr[\mathcal{M}_{\exp}(x)=y]=\frac{\exp\left(\frac{\varepsilon d(x,y)}{2\Delta d}\right)}{\sum_{y^{\prime}\in\mathcal{Y}}\exp\left(\frac{\varepsilon d(x,y^{\prime})}{2\Delta d}\right)},$
where $\Delta d=\max_{y,y^{\prime}\in\mathcal{Y}}|d(x,y)-d(x,y^{\prime})|$ is the sensitivity of the scoring function $d$.
In the context of trajectory collection, the most used scoring function is the negative Euclidean distance $d(x,y)\coloneq-||x-y||_{2}$. Therefore, locations with a smaller Euclidean distance to the private location $x$ are more likely to be sampled.
The biggest weakness of the Exponential mechanism is the complexity in calculating the score function and sampling from the probability distribution. It requires computing the Euclidean distance between every pair of locations in the location space, which is computationally expensive for large-scale spaces.  In ATP and NGram, the location space for perturbation changes at each time step, meaning that the score functions need to be recomputed at each perturbation, resulting in $\mathcal{O}(m)$ time complexity for each perturbation, where $m$ is the size of the location space.
The complexity of sampling from the Exponential mechanism is also $\mathcal{O}(m)$, as it needs to sample from an $m$-piece cumulative distribution function (CDF) at each perturbation.
In contrast, we will see that piecewise-based perturbation mechanisms in TraCS do not require score functions, and they sample from a $3$-piece CDF at each perturbation, resulting in a negligible constant time complexity, i.e. $\Theta(1)$, for each perturbation.
B.2. Dominant Sector Comparison Between the Strawman Approach and TraCS-D
For direction perturbation, a small-size and high-probability dominant sector is desirable for more accurate perturbation. However, these two properties are generally conflicting with each other: a smaller dominant sector typically has a lower probability.  We will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value. Figure 11(b) shows two exemplary quantitative comparisons between them.
Small $\varepsilon$ region. In this case, the dominant sector of TraCS-D is larger than the strawman approach, but the probability of the dominant sector is significantly higher, especially when $k$ is large.
Large $\varepsilon$ region. In such cases, the dominant sector of TraCS-D is significantly smaller than the strawman approach, while maintaining almost the same probability as the strawman approach.
We can observe the effect of $k$ on the strawman approach. Large $k$ values reduce the size difference of the dominant sector size compared with TraCS-D, but they also dramatically reduce the probability of the dominant sector in the small $\varepsilon$ region. Conversely, small $k$ values increase the probability of the dominant sector in the small $\varepsilon$ region, but they also lead to a larger dominant sector in the large $\varepsilon$ region.
(a) $k=6$
(b) $k=12$
Figure 11. Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach.
[…]
Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach.
B.3. Detailed Expression of $|\mathcal{D}_{r(\varphi)}|$
For each location $\tau_{i}$, denote its directions to the four endpoints of the rectangular location space as $\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4}$. We have:
$\begin{split}\varphi_{1}&=\mathrm{atan2}\left(b_{\text{end}}-b_{i},a_{\text{end}}-a_{i}\right),\\ \varphi_{2}&=\mathrm{atan2}\left(b_{\text{end}}-b_{i},a_{\text{sta}}-a_{i}\right),\\ \varphi_{3}&=\mathrm{atan2}\left(b_{\text{sta}}-b_{i},a_{\text{sta}}-a_{i}\right)+2\pi,\\ \varphi_{4}&=\mathrm{atan2}\left(b_{\text{sta}}-b_{i},a_{\text{end}}-a_{i}\right)+2\pi.\end{split}$
Then, $|\mathcal{D}_{r(\varphi)}|$ has the following four cases:
$\begin{split}|\mathcal{D}_{r(\varphi)}|=\begin{dcases}\frac{a_{\text{end}}-a_{i}}{\cos\varphi}&\text{if}\ \varphi\in[0,\varphi_{1})\cup[\varphi_{4},2\pi),\\ \frac{b_{\text{end}}-b_{i}}{\sin\varphi}&\text{if}\ \varphi\in[\varphi_{1},\varphi_{2}),\\ \frac{a_{i}-a_{\text{sta}}}{-\cos\varphi}&\text{if}\ \varphi\in[\varphi_{2},\varphi_{3}),\\ \frac{b_{i}-b_{\text{sta}}}{-\sin\varphi}&\text{if}\ \varphi\in[\varphi_{3},\varphi_{4}),\end{dcases}\end{split}$
depending on the direction $\varphi$ from reference location $\tau_{i}=(a_{i},b_{i})$.
B.4. Redesigned SW
We redesign the SW mechanism to achieve $\varepsilon$-LDP for the circular space $[0,2\pi)$ and the linear space $[0,1)$. The key idea is to transform the original sampling distribution to these spaces while preserving the LDP constraint.  Similar to Definition 3.1, the redesigned SW mechanism for direction perturbation is defined as follows.
Definition B.2 (Redesigned SW for direction perturbation).
Given a private direction $\varphi$ and a privacy parameter $\varepsilon$, redesigned SW for direction perturbation is a mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ defined by:
$pdf[\mathcal{M}_{\circ}(\varphi)=\varphi^{\prime}]=\begin{dcases}p_{\varepsilon}&\text{if}\ \varphi^{\prime}\in[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon}),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=\frac{1}{2\pi\varepsilon}(\exp(\varepsilon)-1)$ is the sampling probability, and $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
$\begin{split}l_{\varphi,\varepsilon}&=\left(\varphi-\pi\frac{\exp(\varepsilon)(\varepsilon-1)+1}{(\exp(\varepsilon)-1)^{2}}\right)\mod 2\pi,\\ r_{\varphi,\varepsilon}&=\left(\varphi+\pi\frac{\exp(\varepsilon)(\varepsilon-1)+1}{(\exp(\varepsilon)-1)^{2}}\right)\mod 2\pi.\end{split}$
Similar to Definition 3.3, the redesigned SW mechanism for distance perturbation is defined as follows.
Definition B.3 (Redesigned SW for distance perturbation).
Given a private distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, redesigned SW for distance perturbation is a mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ that
$pdf[\mathcal{M}_{-}(\overline{r}(\varphi))=\overline{r}^{\prime}(\varphi)]=\begin{dcases}p_{\varepsilon}&\text{if}\ \overline{r}^{\prime}(\varphi)\in[u,v),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=(\exp(\varepsilon)-1)/\varepsilon$ is the sampling probability, and $[u,v)$ is the sampling interval that
$\begin{split}[u,v)&=\begin{cases}\overline{r}(\varphi)+[-C,C)&\text{if}\ \overline{r}^{\prime}(\varphi)\in[C,1-C),\cr[0,2C)&\text{if}\ \overline{r}^{\prime}(\varphi)\in[0,C),\cr[1-2C,1)&\text{otherwise},\end{cases}\end{split}$
with $C=(\exp(\varepsilon)(\varepsilon-1)+1)/(2(\exp(\varepsilon)-1)^{2})$.
Compared with Definition 3.1 and Definition 3.3, the redesigned SW mechanisms have a larger $p_{\varepsilon}$ and a narrower $[l,r)$ or $[u,v]$ interval for sampling. We can treat the redesigned SW mechanisms as a more aggressive perturbation mechanism: it tries to sample from a narrow dominant sector with a higher probability ($p_{\varepsilon}$), which also leads to a higher probability ($p_{\varepsilon}/\exp(\varepsilon)$) of non-dominant sectors due to the LDP constraint.