### v2

# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

Abstract
Trajectory collection is fundamental for location-based services but often involves sensitive information, such as users’ daily activities, raising significant privacy concerns. Local Differential Privacy (LDP) provides strong privacy guarantees for users, even when the data collector is untrusted. Existing trajectory collection methods under LDP are limited to discrete location spaces, where the number of locations affects both privacy guarantees and trajectory utility. Moreover, many real-world scenarios, such as flying trajectories or sensor trajectories of wearable devices, operate in continuous location spaces, making existing methods inadequate.
This paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. We propose two novel methods: TraCS-D, which perturbs the direction and distance of locations, and TraCS-C, which perturbs the Cartesian coordinates of locations. Both methods are theoretically and experimentally analyzed for trajectory utility in continuous spaces. TraCS can also be applied to discrete spaces by rounding perturbed locations to the nearest discrete points. In this case, TraCS’s privacy and utility guarantees are independent of the number of locations in the space, and has only $\Theta(1)$ time complexity in each perturbation generation. Evaluation results on discrete location spaces validate the efficiency advantage and show that TraCS outperforms state-of-the-art methods with improved trajectory utility, especially for large privacy parameters.
1 Introduction
Trajectory data from users—sequences of locations describing movement over time—are a fundamental resource for activity analysis and location-based services, including activity classification and routine detection. However, directly collecting trajectory data raises significant privacy concerns, especially given users’ growing privacy expectations and stringent regulations [1, 19]. Anonymization techniques like $k$-anonymity [28, 20] have proven insufficient to protect trajectory privacy [6]. Thus, stronger privacy guarantees are necessary for such data.
Local differential privacy (LDP) provides a provable worst-case privacy guarantee by ensuring the hardness of distinguishing a user’s data from any released data. The privacy level of an LDP mechanism is quantified by its privacy parameter $\varepsilon$, which also generally implies the extent of data perturbation. Through LDP, users can release their perturbed trajectory data to an untrusted data collector while still benefiting from the services provided by the collector.
Existing LDP methods for trajectory collection are designed for discrete spaces. They leverage discrete LDP mechanisms, such as the Exponential mechanism [27], to perturb the trajectory data. Discrete LDP mechanisms are explicitly defined on the size of the location space. As a result, these methods either rely on properly splitting the entire location space into grids [29] or assume that the location space consists of a finite set of labeled locations (i.e. points of interest) [9, 37]. However, relying on discrete spaces has three issues: (i) Their privacy guarantees are inherently limited to the chosen discrete sets. For instance, a discrete location space with $10$ points always provides a weaker privacy guarantee than one with $100$ points, since any inference attack has at least a $10\%$ chance of success in the former, regardless of the privacy parameter $\varepsilon$. (ii) Their efficacy and efficiency are often constrained by the size of the discrete set. As the size increases, the probability of selecting the true location decreases. Meanwhile, the widely used Exponential mechanism incurs linear sampling complexity when generating each perturbed location, making it inefficient. (iii) Discrete methods are not directly applicable to naturally continuous location spaces, such as flying and sailing trajectories, sensor trajectories from wearable devices, and similar scenarios.
To address these limitations, this paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. A continuous space models locations as continuous values, such as GPS coordinates in $180,180]\times[-90,90]$. Unlike discrete spaces, it contains infinitely many candidate locations. Collecting trajectory data directly in continuous spaces is natural for many applications, and offers three key advantages over discretizing the space and then applying discrete methods: (i) The privacy guarantee holds for the entire continuous space, regardless of the “number” of locations. (ii) The sampling mechanism operates directly on the continuous space, so its efficacy and efficiency do not relate to the “number” of locations. (iii) Furthermore, perturbed locations in the continuous space can be post-processed by rounding to any discrete space within it, making the approach applicable to both continuous and discrete spaces.
Designing trajectory collection methods for continuous spaces under LDP presents unique challenges: they must ensure LDP for every location in the continuous space; they must preserve sufficient trajectory utility for downstream tasks compared to the original trajectories (formalized in Section 4); and they should efficiently generate perturbed locations to support real-time applications.
This paper proposes two LDP methods for trajectory collection in continuous spaces. The main insight is to decompose the 22D continuous space into 11D subspaces and design mechanisms for each subspace. Our mechanisms leverage the concept of piecewise-based mechanisms [31, 21] for 11D bounded numerical domains. [^1] Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms, e.g. the truncated Laplace mechanism [22, 15]. While they can be incorporated into TraCS, they are more complex and less effective than piecewise-based mechanisms. Section 3.5.4 provides details. Based on two different decompositions of the continuous space, we propose two methods: TraCS-D and TraCS-C. The method TraCS-D treats the continuous space as the composition of a direction space and a distance space. We exploit the direction information of trajectories and design a novel direction perturbation mechanism to ensure LDP for the direction space. Combined with a redesigned piecewise-based mechanism for the distance space, TraCS-D ensures LDP for the continuous 22D space. The method TraCS-C treats the continuous space as the Cartesian product of two distance spaces. Each location is represented by a 22D Cartesian coordinate, and TraCS-C perturbs these coordinates. Both methods guarantee LDP for the entire 22D continuous space. They can also be applied to any discrete space within the continuous space by rounding the perturbed locations to the nearest discrete points. Compared with existing methods based on the Exponential mechanism [9, 37], our methods are unaffected by the number of locations in the location space and require only $\Theta(1)$ time to generate each perturbed location. Specifically, our contributions are as follows:
  • To the best of our knowledge, this is the first work to provide trajectory collection methods for continuous space under pure LDP. We propose TraCS-D and TraCS-C for this purpose. These methods introduce novel location perturbation mechanisms that exploit the direction information and coordinate information of trajectories in continuous spaces. We theoretically and experimentally analyze their trajectory utility.
  • Our methods TraCS can also be applied to discrete spaces. Compared with existing methods for discrete spaces, TraCS has significantly lower computation complexity in generating perturbed locations. Experimental results on discrete location spaces validate the efficiency of TraCS, showing an average time cost of less than $1\%$ of NGram, L-SRR, and ATP [9, 37, 29], three state-of-the-art LDP methods designed for discrete spaces. Additionally, the results demonstrate that TraCS generally achieve better trajectory utility, particularly for larger privacy parameters.
Structure. The rest of the paper is organized as follows. Section 2 introduces the preliminaries. Section 3 presents TraCS-D and TraCS-C methods, their theoretical analysis, along with discussions on their differences and extensions. Following that, Section 4 presents the experimental evaluations on continuous and discrete spaces.
2 Preliminaries
This section formulates the problem, introduces LDP, and reviews three trajectory collection methods under LDP, along with their shortcomings in continuous spaces, which motivates the necessity of designing new LDP methods for trajectory collection in continuous spaces.
2.1 Problem Formulation
A typical trajectory collection schema consists of a set of trajectories from users and one collector. Each trajectory $\mathcal{T}$ is a sequence of locations $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, where $\tau_{i}\in\mathcal{S}$ and $\mathcal{S}\subset\mathbb{R}^{2}$ is a continuous and bounded domain. The collector needs to collect these trajectories to provide analysis or location-based services.
However, the collector is untrusted and may act as an adversary attempting to infer users’ sensitive data. Therefore, releasing the sensitive trajectories to the collector poses a privacy risk. To protect privacy, users perturb their trajectories using a privacy mechanism $\mathcal{M}:\mathcal{S}\to Range(\mathcal{M})$, and then send the perturbed trajectories $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$ to the collector.
We aim to design a mechanism $\mathcal{M}$ that satisfies $\varepsilon$-LDP, providing provable privacy guarantee for sensitive trajectories. Additionally, $\mathcal{M}$ should preserve utility, ensuring perturbed trajectories remain comparable to the originals, while incurring only modest computational overhead suitable for real-time applications.
2.2 Local Differential Privacy
Definition 1 ($\varepsilon$-LDP [11]).
A perturbation mechanism $\mathcal{M}:\mathcal{X}\to Range(\mathcal{M})$ satisfies $\varepsilon$-LDP, if for two arbitrary input $x_{1}$ and $x_{2}$, the probability ratio of outputting the same $y$ is bounded:
$\forall x_{1},x_{2}\in\mathcal{X},\forall y\in Range(\mathcal{M}):\frac{\Pr[\mathcal{M}(x_{1})=y]}{\Pr[\mathcal{M}(x_{2})=y]}\leq\exp(\varepsilon).$
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function (pdfpdf). Intuitively, Definition 1 represents the difficulty of distinguishing $x_{1}$ and $x_{2}$ given $y$. Lower values of the privacy parameter $\varepsilon\in[0,+\infty)$ mean higher privacy. For example, $\varepsilon=0$ requires $\mathcal{M}$ to map two arbitrary inputs to any output $y$ with the same probability, thus the perturbed output contains no distribution information of the sensitive input, making any hypothesis-testing method to infer the sensitive $x$ powerless.
Theorem 1 (Sequential Composition of LDP [13, 37]).
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two mechanisms that satisfy $\varepsilon_{1}$ and $\varepsilon_{2}$-LDP, respectively. Their composition, defined as $\mathcal{M}_{1,2}:(\mathcal{X}_{1},\mathcal{X}_{2})\to(Range(\mathcal{M}_{1}),Range(\mathcal{M}_{2}))$, satisfies $(\varepsilon_{1}+\varepsilon_{2})$-LDP.
Sequential composition allows us to design LDP mechanisms for 22D data by composing two mechanisms for 11D.
2.2.1 Piecewise-based Mechanism
If $\mathcal{X}$ is continuous and bounded, the state-of-the-art LDP mechanisms are piecewise-based [31, 21], i.e. they sample the output from piecewise probability distributions defined on $\mathcal{X}$. We summarize them in the following form:
Definition 2 (Piecewise-based mechanism).
A piecewise-based mechanism $\mathcal{M}:\mathcal{X}\to Range(\mathcal{M})$ is a family of probability distributions that, given input $x\in\mathcal{X}$, outputs $y\in Range(\mathcal{M})$ according to
$pdf[\mathcal{M}(x)=y]=\begin{dcases}p_{\varepsilon}&\text{if}\ y\in[l_{x,\varepsilon},r_{x,\varepsilon}),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}$ is the sampling probability w.r.t. $\varepsilon$, $l_{x,\varepsilon},r_{x,\varepsilon})$ is the sampling interval w.r.t. $x$ and $\varepsilon$. $Range(\mathcal{M})\supseteq\mathcal{X}$ is also a continuous and bounded domain.
Piecewise-based mechanisms sample the output $y$ with a high probability $p_{\varepsilon}$ within $l_{x,\varepsilon},r_{x,\varepsilon})$ and a lower probability in the remaining two pieces, while satisfying the $\varepsilon$-LDP constraint. Typical instantiations include PM [31] for mean estimation and SW [21] for distribution estimation. Our perturbation mechanisms also leverage the piecewise-based design but with different instantiations tailored for different data domains.
2.2.2 $k$-$R$ and Exponential Mechanism
If $\mathcal{X}$ is discrete, especially when $|\mathcal{X}|$ is not large [32], a classical LDP mechanism is $k$-$R$ [18].
Definition 3.
$k$-$R$ is a sampling mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{X}$ that, given $|\mathcal{X}|=k$ and input $x\in\mathcal{X}$, outputs $y\in\mathcal{X}$ according to
$\Pr[\mathcal{M}(x)=y]=\begin{dcases}\frac{\exp(\varepsilon)}{|\mathcal{X}|-1+\exp(\varepsilon)}&\text{if}\ y=x,\\ \frac{1}{|\mathcal{X}|-1+\exp(\varepsilon)}&\text{otherwise}.\end{dcases}$
$k$-$R$ outputs the truth $x$ with a higher probability or outputs other values with a lower probability, while satisfying $\varepsilon$-LDP.
When there is a semantic distance (score) function between elements in $\mathcal{X}$, the Exponential mechanism [27] is more widely used. Unlike $k$-$R$, which treats other values (except $x$) equally, the Exponential mechanism leverages a score function to assign different probabilities. $k$-$R$ is a special case of the Exponential mechanism when the score function is a binary indicator function.
Definition 4 (Exponential Mechanism [27]).
Given a score function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, a privacy parameter $\varepsilon>0$, and a set of possible outputs $\mathcal{Y}$, the Exponential mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is defined by:
$\Pr[\mathcal{M}(x)=y]=\frac{\exp\left(\frac{\varepsilon d(x,y)}{2\Delta d}\right)}{\sum_{y^{\prime}\in\mathcal{Y}}\exp\left(\frac{\varepsilon d(x,y^{\prime})}{2\Delta d}\right)},$
where $\Delta d=\max_{x,y,y^{\prime}\in\mathcal{Y}}|d(x,y)-d(x,y^{\prime})|$ is the sensitivity of the score function $d$.
[…]
2.3 Existing Methods
Table 1: Comparison with existing methods.
Main technique LDP guarantee
NGram [9] Hierarchical decomposition Discrete space
L-SRR [29] Staircase $R$ mechanism
ATP [37] Direction perturbation
This paper (TraCS)
Direction* &
coordinate perturbation
Continuous space
  • * designed for continuous direction space (different from ATP).
Table 1 summarizes three state-of-the-art trajectory collection methods under pure LDP and our methods.
NGram [9] hierarchically decomposes the physical space $\mathcal{S}$ into fine-grained discrete regions with semantics. Specifically, $\mathcal{S}$ is first decomposed into small regions $R_{s}$, which are further decomposed by categories into $R_{c}$, and then attached with time information $R_{t}$. A decomposed “gram” might look like $r_{sct}=\{\text{mountain, church, 3am}\}$. This decomposition allows trajectory $\mathcal{T}$ to be represented at a knowledge level, enabling the usage of public knowledge to define reachability, i.e. to constrain possible perturbed trajectories. For example, the above gram $r_{sct}$ is unlikely to be visited at 33am as a church is closed at that time, reducing the possible $\mathcal{T}^{\prime}$. Finally, NGram applies the Exponential mechanism to sample a perturbed trajectory from the reachable set.
L-SRR [29] introduces the staircase randomized response (SRR) mechanism for location perturbation. Unlike $k$-$R$, which assigns only two probabilities—one for the true location and another for all other locations—the SRR mechanism employs a staircase-shaped probability distribution. L-SRR recursively partitions the physical space $\mathcal{S}$ and forms location groups $G_{1},G_{2},\dots,G_{g}$ based on their distances to the true location. By integrating the SRR mechanism, L-SRR assigns higher probabilities to locations in groups that are closer to the true location, and lower probabilities to those farther away. This approach is similar to the Exponential mechanism, but uses group-level distance instead of location-level distance.
ATP [37] uses direction perturbation to limit the set of possible locations. It assumes that the location space is a finite set of labeled locations and represents a trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$ as a sequence of these locations. When perturbing a location $\tau_{i+1}$, ATP first perturbs the direction from $\tau_{i}$ to $\tau_{i+1}$. It divides the direction space into $k$ sectors and uses $k$-$R$ to sample a direction sector. Then, the perturbed location $\tau_{i+1}^{\prime}$ is constrained to be within the sampled sector, thereby limiting the possible $\tau_{i+1}^{\prime}$, which is then sampled using the Exponential mechanism.
We summarize their limitations in the context of continuous spaces as follows:
  • These methods are limited to handling discrete location spaces. The number of locations is a crucial parameter that determines the perturbation mechanism and affects the privacy guarantee. Although it is possible to discretize continuous spaces, proper discretization is challenging to balance the trajectory utility and privacy.
  • Beyond discretization challenges, approaches that rely on the Exponential mechanism (e.g. NGram and ATP) incur high computational cost to generate perturbations and tend to degrade trajectory utility. The other method L-SRR requires forming groups for every location in the space, which leads to the same scalability issue. Additionally, designing effective public knowledge (e.g. NGram) is challenging, particularly in continuous spaces where the location set is infinite and semantic information is difficult to define. Meanwhile, public knowledge is often dataset-specific and also known to the adversary, thereby undermining the privacy guarantee.
Appendix C.1 details the weaknesses of the Exponential mechanism. In summary, before perturbing a location, it requires calculating the score functions $d(x,y)$ of all possible pairs of locations related to $x$. This costs $\Theta(m)$ time, where $m$ is the number of locations in the discrete space. Additionally, each sampling (generating each perturbed location) from the $m$-piece cumulative distribution function (CDF) of the Exponential mechanism also costs $\Theta(m)$ time. This time complexity is not affordable for large-scale location spaces. Moreover, since the sampling probability is determined by the score function, the spatial distribution and relative distances among locations further affect the probability assigned to each location, potentially resulting in utility loss.
To address these limitations, we propose TraCS, a novel trajectory collection method that provides full LDP guarantees for trajectories in continuous spaces. TraCS is specifically designed to overcome the privacy, utility, and efficiency challenges posed by discretization.
3 Our Method: TraCS
This section presents our methods, TraCS-D and TraCS-C, along with their theoretical analyses and extensions. Notations used in this section can be found in Table 3.
3.1 Location Space
We consider the location space constrained by a pair of longitude and latitude lines, i.e. $\mathcal{S}\subset\mathbb{R}^{2}$ is a rectangular area $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, where $a_{\text{sta}},a_{\text{end}}$ denote the longitudes and $b_{\text{sta}},b_{\text{sta}}$ denote the latitudes. Consequently, each location $\tau_{i}\in\mathcal{S}$ has a natural representation as a pair of coordinates $(a_{i},b_{i})$. This representation aligns with real-world location data (e.g. GPS data) and existing trajectory collection methods for discrete spaces [9, 10, 37, 16, 29].
The core idea behind having various LDP methods for the location space $\mathcal{S}$ stems from its multiple decompositions. One decomposition means $\mathcal{S}$ can be represented by several subspaces. For example, in addition to the aforementioned longitude-latitude coordinate representation, each location $\tau_{i}\in\mathcal{S}$ can also be represented by a direction component and a distance component. Based on these decompositions, we propose two LDP methods: TraCS-D, which perturbs the direction and distance subspaces, and TraCS-C, which perturbs subspaces of the Cartesian coordinates.
3.2 TraCS-D
Continuous space $\mathcal{S}$ can be represented as the composition of a direction space $\mathcal{D}_{\varphi}$ and a distance space $\mathcal{D}_{r(\varphi)}$ at each (reference) location $\tau_{i}$. Figure 1(a) illustrates this decomposition.
(a) TraCS-D: $\mathcal{S}=\mathcal{D}_{\varphi}\otimes\mathcal{D}_{r(\varphi)}$
(b) TraCS-C: $\mathcal{S}=\mathcal{D}_{a}\times\mathcal{D}_{b}$
Figure 1: Two decompositions of location space $\mathcal{S}$ at $\tau_{i}$. Any other location $\tau_{j}\in\mathcal{S}$ can be represented by direction-distance coordinates $(\varphi,r(\varphi))$ or Cartesian coordinates $(a,b)$.
Direction space $\mathcal{D}_{\varphi}$. In the space $\mathcal{S}\subset$ $\mathbb{R}^{2}$, the direction is represented by the angle $\varphi$ relative to a reference direction. We define the reference direction as the latitude lines, i.e. the $0^{\circ}$ direction. Thus, the direction space $\mathcal{D}_{\varphi}$ is a circular domain $0,2\pi)$ for any location. Fixing a location $\tau_{i}$, any other location $\tau_{j}\neq\tau_{i}$ has a unique direction with respect to $\tau_{i}$.
Distance space $\mathcal{D}_{r(\varphi)}$. The other subspace, i.e. the distance space $\mathcal{D}_{r(\varphi)}$, is determined by the direction $\varphi$. Its size is the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ in the direction $\varphi$, which is a function of $\varphi$. Since $\mathcal{S}$ is rectangular, $r(\varphi)$ has a closed-form expression, which is detailed later. Thus, fixing $\tau_{i}$ and a direction $\varphi$, any other location $\tau_{j}$ has a unique distance $r(\varphi)$ with respect to $\tau_{i}$.
By this decomposition, we can represent any location $\tau_{j}\in S$ as coordinates $(\varphi,r(\varphi))\in(\mathcal{D}_{\varphi},\mathcal{D}_{r(\varphi)})$ relative to $\tau_{i}$. As an example, $\tau_{j}$ in Figure 1(a) is represented by $(\varphi_{1},r(\varphi_{1}))$. This representation is unique and reversible: any pair of coordinates $(\varphi,r(\varphi))$ can be mapped back to a unique location $\tau_{j}\in\mathcal{S}$. Thus, if $\tau_{j}$ is a sensitive location, an LDP mechanism applied to $\tau_{j}=(\varphi,r(\varphi))$ is essentially applied to $\varphi$ and $r(\varphi)$. By Composition Theorem 1, we can design perturbation mechanisms for $\varphi$ and $r(\varphi)$, respectively, to achieve LDP for $\tau_{j}$.
(a) $\varepsilon=4$
(b) $\varepsilon=6$
Figure 2: Comparison of the dominant sectors of the strawman approach and TraCS-D with $\varepsilon=4$ and $\varepsilon=6$. The red angular arcs indicate the dominant sectors, with their probabilities shown on the right. The dominant sector of TraCS-D narrows as $\varepsilon$ increases, leading to a smaller inner-sector (sampling) error. In contrast, the strawman approach has a fixed dominant sector, which is independent of $\varepsilon$ and leads to a large inner-sector error.
3.2.1 Direction Perturbation
The technical challenge in direction perturbation is designing an LDP mechanism for the circular domain $\mathcal{D}_{\varphi}=[0,2\pi)$. This is non-trivial due to the following properties of the circular domain:
  • The circular domain $0,2\pi)$ is bounded but not finite, making the unbounded mechanisms and discrete mechanisms inapplicable.
  • The circular domain has a unique distance metric, e.g. the distance between 0 and $2\pi$ is 0, not $2\pi$, making the Euclidean distance-based mechanisms inapplicable.
Specifically, classical mechanisms like Laplace and Gaussian [13] add unbounded noise, resulting in an output domain of $(-\infty,+\infty)$. This output domain makes them inapplicable to the circular domain $0,2\pi)$. Additionally, these mechanisms (also existing piecewise-based mechanisms) are designed for the distance metric d(y,x)=|yx|d(y,x)=|y-x|, which is inconsistent with the distance in circular domain. Discrete mechanisms, such as the $k$-$R$ mechanism [18], require a finite domain size $k$, making them unsuitable for direct application to the continuous circular domain $0,2\pi)$.
To highlight the necessity and superiority of TraCS-D’s piecewise-based design, we start by presenting a strawman approach that extends the $k$-$R$ mechanism to the circular domain $0,2\pi)$.
Strawman approach: $k$-$R$ + uniform sampling. A former work ATP [37] divides the direction space $0,2\pi)$ into $k$ sectors and applies the $k$-$R$ mechanism. This approach ensures LDP for the $k$ sectors but not for the entire $0,2\pi)$ domain. Even so, we can extend this approach to $0,2\pi)$ by further uniformly sampling a direction $\varphi^{\prime}$ from the output sector of $k$-$R$.
Specifically, the circular domain $0,2\pi)$ is divided into $k$ sectors $i-1,i)\cdot 2\pi/k$ for $i=1,2,\ldots,k$. Assume the sensitive direction $\varphi$ falls into the $j$-th sector. Applying the $k$-$R$ mechanism outputs sector $i=j$ with probability $p=\exp(\varepsilon)/(k-1+\exp(\varepsilon))$, or outputs sector $i\neq j$ with probability $p/\exp(\varepsilon)$. Then, we uniformly sample a direction $\varphi^{\prime}$ from the $i$-th sector, i.e. $\varphi^{\prime}\sim U(i-1,i)\cdot 2\pi/k$.
Limitation. This strawman approach introduces inherent inner-sector errors due to uniform sampling within a specific sector. That is, even though the perturbed direction $\varphi^{\prime}$ falls into the same sector as the sensitive direction $\varphi$, the distance between $\varphi$ and $\varphi^{\prime}$ may still be large due to the large sector size. This issue is particularly prominent when $k$ is small, as each sector then spans $2\pi/k$ radians. Uniform sampling within such wide sectors introduces substantial inner-sector errors, which persist regardless of how large the privacy parameter $\varepsilon$ is set. On the other hand, increasing $k$ to reduce sector width impairs the utility of the $k$-$R$ mechanism [32].
Design rationale. To address this limitation, we design a direction perturbation mechanism in which the perturbed direction is independent of hyperparameters like $k$, thereby avoiding inherent errors beyond those introduced by the privacy parameter $\varepsilon$.
We leverage the idea of piecewise-based mechanisms’ design. Specifically, (i) we instantiate a piecewise-based mechanism for the circular domain $0,2\pi)$, which guarantees LDP for the entire direction space; (ii) the perturbed direction is sampled from a piecewise probability distribution, which is centered around the sensitive direction with high probability and depends solely on the privacy parameter $\varepsilon$, thus eliminating the need for fixed sectors.
Definition 5 (Direction Perturbation Mechanism).
Given a sensitive direction $\varphi$ and a privacy parameter $\varepsilon$, TraCS-D’s direction perturbation mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ is defined by:
$pdf[\mathcal{M}_{\circ}(\varphi)=\varphi^{\prime}]=\begin{dcases}p_{\varepsilon}&\text{if}\ \varphi^{\prime}\in[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon}),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=\frac{1}{2\pi}\exp(\varepsilon/2)$ is the sampling probability, and $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
$\begin{split}l_{\varphi,\varepsilon}&=\left(\varphi-\pi\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}\right)\mod 2\pi,\\ r_{\varphi,\varepsilon}&=\left(\varphi+\pi\frac{\exp(\varepsilon/2)-1}{\exp(\varepsilon)-1}\right)\mod 2\pi.\end{split}$
The above mechanism $\mathcal{M}_{\circ}$ is defined by a piecewise probability distribution with three pieces, where the central piece $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ has a higher probability density $p_{\varepsilon}$. As a piecewise-based mechanism, it evidently satisfies LDP for the circular domain $0,2\pi)$; see Appendix B.1 for the proof.
A key advantage of TraCS-D’s direction perturbation mechanism is the “dominant” sector. We refer to the central piece $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ as the dominant sector because it centers around the sensitive direction $\varphi$ with high probability. The dominant sector adapts dynamically to the privacy parameter $\varepsilon$: as $\varepsilon$ increases, the dominant sector becomes narrower and more tightly centered around the sensitive direction $\varphi$. This adaptivity effectively mitigates the inner-sector errors inherent in the strawman approach, where the sector width is fixed and independent of $\varepsilon$.
Example 1.
Figure 2 compares the strawman approach and TraCS-D. Assume the next location $\tau_{i+1}$ is a sensitive location needing direction perturbation, thus the sensitive direction is $\varphi:\tau_{i}^{\prime}\to\tau_{i+1}$. The dominant sector of TraCS-D is much smaller as $\varepsilon$ increases. For instance, when $\varepsilon=6$ and $\varphi=\pi/6$, it can be calculated that $p_{\varepsilon}\approx 3.20$ and $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})\approx[0.119\pi,0.214\pi)$. Thus, the dominant sector has a size of $0.095\pi\approx 17^{\circ}$ with probability $0.095\pi\times 3.20\approx 0.95$ of being chosen. In comparison, the strawman approach has a dominant sector of $2\pi/6=60^{\circ}$ with probability $0.98$ of being chosen. With almost the same probability of being chosen, TraCS-D’s dominant sector is significantly narrower than that of the strawman approach, resulting in a much smaller inner-sector (sampling) error.
Detailed comparison between the strawman approach and Definition 5 is provided in Appendix C.2. This Appendix will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value.
3.2.2 Distance Perturbation
The other subspace of TraCS-D is the distance space $\mathcal{D}_{r(\varphi)}$. Besides the boundedness, the size of this space also varies with the reference location $\tau_{i}$ and the direction $\varphi$. The technical challenge in designing an LDP mechanism for this space lies in calculating the size $|\mathcal{D}_{r(\varphi)}|$ for any $\varphi$ at any $\tau_{i}$.
Given location $\tau_{i}$, any other location $\tau_{j}\in\mathcal{S}$ can be represented by a pair of coordinates $(\varphi,r(\varphi))$ relative to $\tau_{i}$. Since we focus on $\mathcal{S}$ as a rectangular area $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$, the distance from $\tau_{i}$ to the boundary of $\mathcal{S}$ has a closed-form expression depending on the direction $\varphi$. Specifically, this distance falls into one of the following four cases:
$|\mathcal{D}_{r(\varphi)}|\in\left\{\frac{a_{\text{end}}-a_{i}}{\cos(\varphi)},\frac{b_{\text{end}}-b_{i}}{\sin(\varphi)},\frac{a_{i}-a_{\text{sta}}}{-\cos(\varphi)},\frac{b_{i}-b_{\text{sta}}}{-\sin(\varphi)}\right\},$ (1)
depending on the value of $\varphi$. Appendix C.3 provides the detailed equation form of Equation (1). Figure 3 illustrates the first two cases of $|\mathcal{D}_{r(\varphi)}|$: when $\varphi$ falls into $0,\varphi_{1})$, i.e. Figure 3(a), $|\mathcal{D}_{r(\varphi)}|=(a_{\text{end}}-a_{i})/\cos(\varphi)$; Figure 3(b) shows the second case, i.e. when $\varphi\in[\varphi_{1},\varphi_{2})$, then $|\mathcal{D}_{r(\varphi)}|=(b_{\text{end}}-b_{i})/\sin(\varphi)$.
With the known distance space $0,|\mathcal{D}_{r(\varphi)}|]$ and sensitive distance $r(\varphi)$, we can design a distance perturbation mechanism to guarantee LDP for $\mathcal{D}_{r(\varphi)}$. For the convenience of presentation, we normalize the distance $r(\varphi)$ by $|\mathcal{D}_{r(\varphi)}|$, resulting in a normalized distance $\overline{r}(\varphi)\in[0,1)$. [^2] We use $[0,1)$ instead of $[0,1]$ to ease the presentation of Definition 6 and to align with the circular domain. These two domains are equivalent in implementation. We then design a piecewise-based mechanism for the normalized distance $\overline{r}(\varphi)$ to guarantee LDP on $[0,1)$.
Definition 6 (Distance Perturbation Mechanism).
Given a sensitive distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, TraCS-D’s distance perturbation mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ is defined by:
$pdf[\mathcal{M}_{-}(\overline{r}(\varphi))=\overline{r}^{\prime}(\varphi)]=\begin{dcases}p_{\varepsilon}&\text{if}\ \overline{r}^{\prime}(\varphi)\in[u,v),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=\exp(\varepsilon/2)$ is the sampling probability, and $[u,v)$ is the sampling interval that
$\begin{split}[u,v)&=\begin{cases}\overline{r}(\varphi)+[-C,C)&\text{if}\ \overline{r}(\varphi)\in[C,1-C),\cr[0,2C)&\text{if}\ \overline{r}(\varphi)\in[0,C),\cr[1-2C,1)&\text{otherwise},\end{cases}\end{split}$
with $C=(\exp(\varepsilon/2)-1)/(2\exp(\varepsilon)-2)$.
Similar to $\mathcal{M}_{\circ}$, the above mechanism $\mathcal{M}_{-}$ has a higher probability density $p_{\varepsilon}$ in the central piece $[u,v)$, and a larger $\varepsilon$ results in a higher $p_{\varepsilon}$ and a narrower $[u,v)$, boosting the utility. It ensures LDP for $[0,1)$ and outputs a perturbed normalized distance $\overline{r}^{\prime}(\varphi)$. To cooperate with the direction perturbation, $\overline{r}^{\prime}(\varphi)$ can be mapped to another distance space $\mathcal{D}_{r(\varphi^{\prime})}$ at the perturbed direction $\varphi^{\prime}$. This is a linear mapping without randomness, so the post-processing property [13] ensures it preserves the same privacy level.
(a) $\varphi\in[0,\varphi_{1})$
(b) $\varphi\in[\varphi_{1},\varphi_{2})$
Figure 3: Two cases of $|\mathcal{D}_{r(\varphi)}|$ at $\tau_{i}$ (blue lines).
3.2.3 Workflow of TraCS-D
Combining the direction perturbation and distance perturbation, TraCS-D perturbs each sensitive location in the trajectory.
Algorithm 1 presents the workflow of TraCS-D. It takes the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input. Since TraCS-D relies on a non-sensitive reference location, we add a dummy location $\tau_{0}^{\prime}$ to $\mathcal{T}$, which can be the starting coordinate of $\mathcal{S}$ or a randomly drawn location. Then, TraCS-D iterates over each pair of consecutive locations $\tau_{i}$ and $\tau_{i+1}$. For each sensitive location $\tau_{i+1}$, it perturbs the direction (red block) and normalized distance (blue block) with privacy parameters $\varepsilon_{d}$ and $\varepsilon-\varepsilon_{d}$, respectively. The perturbed distance is then mapped along the perturbed direction (line 9). Line 12 updates the reference location to the perturbed location $\tau_{i+1}^{\prime}$ in place, which is then used as the reference for the next iteration ($i+1$). Finally, the algorithm outputs the perturbed trajectory $\mathcal{T}^{\prime}$.
Algorithm 1 uses each $\tau_{i}^{\prime}$ as the reference location for the next perturbation. In fact, it can be any non-sensitive location in $\mathcal{S}$. We choose different $\tau_{i}^{\prime}$ for alleviating the impact of a specific reference location.
Input: Rectangular location space $\mathcal{S}$, sensitive trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$
Output: Perturbed trajectory $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$
$\mathcal{T}^{\prime}\leftarrow\emptyset,\mathcal{T}\leftarrow\tau_{0}^{\prime}\cup\mathcal{T}$;
[…]
1 for $i\leftarrow 0$ to n1n-1 do
[…]
2  $\tau_{i}=(a_{i},b_{i}),\tau_{i+1}=(a_{i+1},b_{i+1})$;
[…]
3  $\varphi\leftarrow\text{atan2}(b_{i+1}-b_{i},a_{i+1}-a_{i})$;
[…]
4  $\varphi^{\prime}\leftarrow\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$;
5  $R\leftarrow|\mathcal{D}_{r(\varphi)}|$ in Equation (1);
  $\overline{r}(\varphi)\leftarrow\frac{a_{i+1}-a_{i}}{R\cos(\varphi)}$;
[…]
  $\overline{r}^{\prime}(\varphi)\leftarrow\mathcal{M}_{-}(\overline{r}(\varphi);\varepsilon-\varepsilon_{d})$;
[…]
  $R^{\prime}\leftarrow|\mathcal{D}_{r(\varphi^{\prime})}|$, $r^{\prime}(\varphi^{\prime})\leftarrow\overline{r}^{\prime}(\varphi)\times R^{\prime}$;
[…]
6  $\tau_{i+1}^{\prime}=(a_{i}+r^{\prime}(\varphi^{\prime})\cos(\varphi^{\prime}),b_{i}+r^{\prime}(\varphi^{\prime})\sin(\varphi^{\prime}))$;
7  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
  $\tau_{i+1}\leftarrow\tau_{i+1}^{\prime}$;
[…]
8
9return $\mathcal{T}^{\prime}$;
Algorithm 1 TraCS-D
3.2.4 Analysis of TraCS-D
This subsection analyzes the privacy, computational complexity, and utility of TraCS-D.
Theorem 2.
TraCS-D (Algorithm 1) satisfies $n\varepsilon$-LDP for the rectangular location space $\mathcal{S}$.
Proof.
(Sketch) The direction perturbation mechanism $\mathcal{M}_{\circ}$ satisfies $\varepsilon_{d}$-LDP by its definition. For the distance perturbation, the randomness comes entirely from mechanism $\mathcal{M}_{-}$, which satisfies $(\varepsilon-\varepsilon_{d})$-LDP. The post-processing property preserves the same privacy level after linearly mapping the perturbed distance. Then each location satisfies $\varepsilon$-LDP by Composition Theorem 1 (or by computing the 2D pdfpdf ratio in the LDP definition). Hence, the entire perturbed trajectory satisfies $n\varepsilon$-LDP. Appendix B.2 provides details. ∎
Complexity. TraCS-D (Algorithm 1) has $\Theta(1)$ time complexity for each location, as each line is $\Theta(1)$. Thus, the entire algorithm has $\Theta(n)$ time complexity, where $n$ is the length of the trajectory. The space complexity is also $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$.
The $\Theta(n)$ time complexity is optimal for per-location perturbation mechanisms. Among existing LDP mechanisms for discrete location spaces, even if we ignore the complexity of the score function and sampling in the Exponential mechanism [27], NGram [9] has $\Theta(tn)$ time complexity, where $t$ is the number of predefined “gram”. It also includes a large constant factor due to its search to satisfy reachability constraints. L-SRR [29] has $\Theta(mn)$ time complexity, where $m$ is the number of locations, due to the grouping of locations. ATP [37] also has $\Theta(mn)$ time complexity, because of its trajectory merging step.
Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can be reduced to $\Theta(1)$ by only updating each perturbed location $\tau_{i+1}^{\prime}$ in place within $\mathcal{T}$. This results in a lightweight memory footprint, which is crucial for edge computing.
Theorem 3.
In TraCS-D, both the worst-case mean square error (MSE) of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ converge to zero with a rate of $\Theta(e^{-\varepsilon/2})$.
Proof.
(Sketch) The MSE can be calculated by the closed-form expression of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$, allowing us to prove the convergence rate. Appendix B.3 provides the details. ∎
This theorem indicates that both the error of the perturbed direction and distance are reduced exponentially with the privacy parameter $\varepsilon$. Note that the size of the distance space also affects the variance. Specifically, denote the output of $\mathcal{M}_{-}$ as a random variable $Y$ and the sensitive input as $x$. After the linear mapping, the perturbed distance is $Y\cdot|\mathcal{D}_{r}|$ and the sensitive distance is $x\cdot|\mathcal{D}_{r}|$, where $|\mathcal{D}_{r}|$ is the size of the distance space. Thus, the MSE of the perturbed distance is $\text{MSE}[Y\cdot|\mathcal{D}_{r}|]=|\mathcal{D}_{r}|^{2}\cdot\text{MSE}[Y]$. Therefore, a smaller $|\mathcal{D}_{r}|$ leads to a smaller MSE of the perturbed distance.
3.3 TraCS-C
TraCS-C decomposes the rectangular location space $\mathcal{S}$ into two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, as shown in Figure 1(b). Specifically, we can fix $(a_{\text{sta}},b_{\text{sta}})$ as the reference location, with longitude $a_{\text{sta}}$ as the $a$-axis and latitude $b_{\text{sta}}$ as the $b$-axis. Then any location $\tau_{i}\in\mathcal{S}$ can be represented as a pair of Cartesian coordinates $(d_{a},d_{b})$, where $d_{a}\in\mathcal{D}_{a}$ and $d_{b}\in\mathcal{D}_{b}$ are the distances from $\tau_{i}$ to the $a$-axis and $b$-axis, respectively. This representation is unique and reversible: any pair of distances $(d_{a},d_{b})$ can be mapped back to a unique location $\tau_{i}\in\mathcal{S}$. Thus, if $\tau_{i}$ is a sensitive location, an LDP mechanism applied to $\tau_{i}$ is essentially applied to $d_{a}$ and $d_{b}$. By Composition Theorem 1, we can achieve LDP for $\tau_{j}$ by applying perturbation mechanisms for $d_{a}$ and $d_{b}$, respectively.
3.3.1 Coordinates Perturbation
In TraCS-C’s representation of locations, $\mathcal{D}_{a}=[0,a_{\text{end}}-a_{\text{sta}})$ and $\mathcal{D}_{b}=[0,b_{\text{end}}-b_{\text{sta}})$ are independent of any specific location. Therefore, the location perturbation of any $\tau_{i}$ can be performed independently in $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$.
Following the same idea of distance perturbation in TraCS-D, we normalize the distance $d_{a}$ and $d_{b}$ to $[0,1)$ by dividing $|\mathcal{D}_{a}|$ and $|\mathcal{D}_{b}|$, respectively. Then the mechanism $\mathcal{M}_{-}$ in Definition 6 provides $\varepsilon$-LDP for the normalized distances, and the linear mappings back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ preserve the privacy guarantee.
Input: Rectangular location space $\mathcal{S}$, sensitive trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$
Output: Perturbed trajectory $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$
1 $\mathcal{T}^{\prime}\leftarrow\emptyset$;
2 for $i\leftarrow 1$ to $n$ do
3  $\tau_{i}=(a_{i},b_{i})$;
[…]
4  $(\overline{d}_{a},\overline{d}_{b})\leftarrow(\frac{a_{i}-a_{\text{sta}}}{a_{\text{end}}-a_{\text{sta}}},\frac{b_{i}-b_{\text{sta}}}{b_{\text{end}}-b_{\text{sta}}})$;
[…]
5  $\overline{d}^{\prime}_{a}\leftarrow\mathcal{M}_{-}(\overline{d}_{a};\varepsilon/2)$;
6  $\overline{d}^{\prime}_{b}\leftarrow\mathcal{M}_{-}(\overline{d}_{b};\varepsilon/2)$;
[…]
7  $(d_{a}^{\prime},d_{b}^{\prime})\leftarrow(\overline{d}^{\prime}_{a}|\mathcal{D}_{a}|,\overline{d}^{\prime}_{b}|\mathcal{D}_{b}|)$, $\tau^{\prime}_{i}=(a_{\text{sta}},b_{\text{sta}})+(d_{a}^{\prime},d_{b}^{\prime})$;
8  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{\tau^{\prime}_{i}\}$;
9
10return $\mathcal{T}^{\prime}$;
Algorithm 2 TraCS-C
Algorithm 2 shows the workflow of TraCS-C. It takes the rectangular location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input, and outputs the perturbed trajectory $\mathcal{T}^{\prime}$. For each location $\tau_{i}=(a_{i},b_{i})$ in $\mathcal{T}$, Line 4 computes its coordinates $(d_{a},d_{b})$ and normalizes them to $(\overline{d}_{a},\overline{d}_{b})\in[0,1)\times[0,1)$. Lines 5-6 perturb the normalized coordinates $\overline{d}_{a}$ and $\overline{d}_{b}$ using $\mathcal{M}_{-}$ with privacy parameter $\varepsilon/2$. Then Line 8 maps the perturbed normalized coordinates back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, and converts them to the longitude-latitude representation $\tau^{\prime}_{i}$.
3.3.2 Analysis of TraCS-C
Privacy and complexity of TraCS-C can be analyzed similarly to TraCS-D.
Theorem 4.
TraCS-C (Algorithm 2) satisfies $n\varepsilon$-LDP for the rectangular location space $\mathcal{S}$.
Proof.
(Sketch) Algorithm 2 uses mechanism $\mathcal{M}_{-}$ twice, each with a privacy parameter $\varepsilon/2$. By Composition Theorem 1 (or by computing the 2D pdfpdf ratio in the LDP definition), their composition satisfies $\varepsilon$-LDP. The subsequent linear mappings to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ are post-processing steps that preserve the same privacy. Hence, each perturbed location satisfies $\varepsilon$-LDP, and the whole trajectory satisfies $n\varepsilon$-LDP. Appendix B.4 provides details. ∎
Complexity. TraCS-C (Algorithm 2) has $\Theta(n)$ time complexity and $\Theta(n)$ space complexity, where $n$ is the length of the trajectory. The time complexity is $\Theta(n)$ because each line of Algorithm 2 is $\Theta(1)$. The space complexity is $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$. Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can also be reduced to $\Theta(1)$ by updating each perturbed location $\tau_{i}^{\prime}$ in place within $\mathcal{T}$, which benefits edge devices with limited computation resources.
3.4 Rounding to Discrete Space
Although TraCS is designed for continuous location spaces, it can be applied to any fine-grained discrete location space by rounding the perturbed locations to the nearest discrete locations. The rounding is a post-processing step that does not affect the privacy guarantee for the continuous space.
Specifically, if the discrete location space is a set of cells, we can round the perturbed location to the cell that contains it. Formally, assume $\mathcal{C}=\{c_{1},\ldots,c_{m}\}\subseteq\mathcal{S}$ discretize $\mathcal{S}$ into $m$ cells. If $\tau_{i}^{\prime}$ is the perturbed location of $\tau_{i}$ in TraCS, then the discretized TraCS outputs $c_{j}$ such that $\tau_{i}^{\prime}\in c_{j}$. Commonly, the cells are evenly divided, so the rounding costs $\Theta(1)$ time, as $c_{j}$ can be determined by the coordinates of $\tau_{i}^{\prime}$.
[…]
Heuristic reference location. The possible locations in a discrete space are generally public knowledge, which aids in heuristically selecting a better reference location for TraCS-D. The key observation is that a smaller distance space $\mathcal{D}_{r(\varphi)}$ leads to a smaller distance perturbation error. Therefore, it is beneficial to select a reference location that minimizes the size of $\mathcal{D}_{r(\varphi)}$ to other locations.
Specifically, the direction perturbation from the reference location is likely to be within the dominant sector shown in Figure 2. If a dominant sector contains most of the locations (resulting in smaller rounding error) and is near the boundary of the location space (resulting in smaller distance space), then the corresponding reference location is a good choice.
3.5 Discussions and Extensions
This subsection discusses the comparison between TraCS-D and TraCS-C, including their privacy, sensitivity to the location space, and extensions to other types of location spaces.
3.5.1 Comparison of Privacy Between TraCS-D and TraCS-C
Although TraCS-D and TraCS-C provide the same level of privacy quantified by $\varepsilon$-LDP—more specifically, location-level (or event-level) LDP as described in surveys [26, 3]—they have different privacy interpretations.
In TraCS-D, the LDP guarantee is provided for the direction information and the subsequent distance information. This means that when the usage of TraCS-D and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive direction $\varphi:\tau^{\prime}_{i}\to\tau_{i+1}$ and the sensitive distance $r(\varphi):|\tau^{\prime}_{i}\to\tau_{i+1}|$ from the known $\tau^{\prime}_{i}$ and $\tau^{\prime}_{i+1}$. Although the distance space $\mathcal{D}_{r(\varphi)}$ relies on a specific direction, it does not leak the direction information, as the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of any location and TraCS-D uses the perturbed direction.
In TraCS-C, the LDP guarantee is provided for the two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$. This means that when the usage of TraCS-C and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive distances $d_{a}$ and $d_{b}$.
3.5.2 Impact of the Size of $\mathcal{S}$
Intuitively, a larger location space $\mathcal{S}$ requires larger privacy parameters to achieve the same level of utility. For TraCS-D and TraCS-C, this impact is reflected in the size of the direction space and distance spaces.
In TraCS-D, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of the size of $\mathcal{S}$, meaning that the MSE of the perturbed direction, $\mathrm{MSE}[\varphi^{\prime}]$, remains constant regardless of the size of $\mathcal{S}$. The distance space $\mathcal{D}_{r(\varphi)}$ depends on the specific direction $\varphi$ and the size of $\mathcal{S}$. In the worst case, $\mathcal{D}_{r(\varphi)}$ is the diagonal of $\mathcal{S}$. We have shown that the MSE of $\mathcal{M}_{-}$ is quadratic to the size of the distance space after linear mapping in Section 3.2.4. Therefore, when $|\mathcal{D}_{r(\varphi)}|$ increases linearly, the expected error of the perturbed distance, i.e. $\mathrm{E}[|r^{\prime}(\varphi)-r(\varphi)|]$, also increases linearly. Nonetheless, the error in the direction space $\mathcal{D}_{\varphi}$ affects the error in $\mathcal{D}_{r(\varphi)}$ in a non-linear way. Specifically, under the Euclidean distance metric, this effect is $|\mathcal{D}_{r(\varphi)}|\cdot 2\sin(|\varphi-\varphi^{\prime}|/2)$, which is the chord length of the arc between $\varphi$ and $\varphi^{\prime}$.
In TraCS-C, both distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ are determined by the size of $\mathcal{S}$. Since they use mechanism $\mathcal{M}_{-}$ independently, their expected error increases linearly with the increase of $|\mathcal{D}_{a}|$ and $|\mathcal{D}_{b}|$. Meanwhile, the error in $\mathcal{D}_{a}$ affects the error in $\mathcal{D}_{b}$ linearly.
3.5.3 Comparison with Geo-indistinguishability
Geo-indistinguishability (Geo-Ind for brevity, also called metric privacy in recent works) [2] is a more general (or weaker) version of LDP. From the perspective of Geo-ind, LDP force the proximity-aware function $d(x,y)$ in Geo-ind to be $d(x,y)=1$ for all $x,y\in S$, ignoring that different $x$ and $y$ have different $d(x,y)$ in the context of their locations. Conversely, if the upper bound $\max_{x,y\in S}d(x,y)\leq d^{*}$ holds in $S$, then TraCS also satisfies $d^{*}\varepsilon$-Geo-ind. However, this trivial satisfaction of $d^{*}\varepsilon$-Geo-ind is highly suboptimal, since it uses the global upper bound of $d(x,y)$ rather than the fine-grained, actual values of $d(x,y)$ for different location pairs of $x$ and $y$.
We omit trajectory utility evaluations against Geo-ind as it is a different privacy model from LDP. Generally, a well-designed Geo-ind mechanism can achieve better data utility than LDP mechanisms. For piecewise-shaped mechanisms, it is possible to tailor different pieces to account for varying $d(x,y)$ for different $x,y$, thereby satisfying Geo-ind while achieving better data utility than the current LDP mechanisms.
3.5.4 Other LDP Mechanisms for Bounded and Continuous Space
In this paper, we leverage piecewise-based mechanisms for bounded numerical domains to design the direction and distance perturbations in TraCS. Other LDP mechanisms for bounded numerical domains can also be incorporated into TraCS, but most of them may require redesign to suit the direction perturbation, as the circular domain $\mathcal{D}_{\varphi}=[0,2\pi)$ has a different distance metric than linear domains. One exception is the Purkayastha mechanism [33], which is designed for sphere $\mathbb{S}^{n-1}$, e.g. circular lines when $n=2$. It can be an alternative to $\mathcal{M}_{\circ}$, but not be directly applicable to $S\subset\mathbb{R}^{2}$. We don’t leverage it for direction perturbation because of its sampling complexity (Section 3.4.2 in [33]) and higher MSE.
Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms, including truncated Laplace [22, 15] and truncated Gaussian mechanisms [4]. These mechanisms modify Laplace and Gaussian distributions by redesigning distributions on a bounded domain, e.g. $[0,1)$. However, truncated mechanisms cannot preserve the same privacy level as their original counterparts [22, 15, 4]. Determining their new privacy level is non-trivial and typically relies on numerical testing algorithms [15, 4] rather than closed-form solutions. Compared with truncated mechanisms, piecewise-based mechanisms are explicitly designed for LDP in bounded numerical domains, making them more effective.
3.5.5 More Extensions
Appendix C.4 provides two more extensions of TraCS and experimental results, including (i) other piecewise-based mechanisms and (ii) other shapes of location space. The experimental evaluations are included in Appendix C.6.
4 Evaluation
This section evaluates the performance of TraCS on both continuous and discrete spaces.
4.1 Results on Continuous Space
The strawman method, our extension of $k$-$R$ for direction perturbation, can be combined with the distance perturbation mechanism in TraCS-D to ensure LDP in continuous spaces $\mathcal{S}\subset\mathbb{R}^{2}$. For the name simplicity, we still refer to this extended method as the strawman method and compare it with TraCS in continuous space.
4.1.1 Setup
TraCS requires the following parameters: the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, the privacy parameter $\varepsilon$, and $\varepsilon_{d}$ for direction perturbation in TraCS-D. The strawman method also requires an instance of $k$ in $k$-$R$. We use configurations:
  • Location space $\mathcal{S}$: We consider synthetic and real-world settings. (i) The default location space is $0,1)\times[0,1)$. Without loss of generality, we consider a larger space $\mathcal{S}=[0,2)\times[0,10)$ for comparison. We generate $100$ random trajectories for each location space as synthetic datasets. This setup excludes dataset-specific effects to emphasize algorithmic performance. (ii) Areas of TKY and CHI: These two trajectory datasets were collected in Tokyo and Chicago, extracted from the Foursquare dataset [34] and Gowalla dataset [7], respectively. Locations in TKY are within area $139.47,139.90)\times[35.51,35.86)$ and CHI within area $87.9,-87.5)\times[41.6,42.0)$. We use the first $100$ trajectories in each dataset.
  • Privacy parameter $\varepsilon_{d}$: We set $\varepsilon_{d}=\varepsilon\pi/(\pi+1)$ to heuristically balance the perturbation in $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ according to their domain sizes.
  • $k$-$R$: We set $k=6$ as the strawman method’s default value, and consider $3$-$R$ and $12$-$R$ for comparison.
We evaluate $\varepsilon$ in the range $[2,10]$. For 2D discrete spaces with large location sets, this range represents a moderate privacy level. [^3] In Definition 1, the privacy parameter $\varepsilon$ is defined with respect to a specific domain $\mathcal{X}$; its practical strength therefore depends on the size of $\mathcal{X}$. For example, with $k$-$R$ mechanism and $|\mathcal{X}|=3600$ locations, $\varepsilon=10$ yields the true location output with probability $\approx 0.85$, while $\varepsilon=2$ results in a much lower probability of $\approx 0.002$. The Exponential mechanism exhibits similar behavior, though the exact probabilities depend on the score function. In such cases where $|\mathcal{X}|$ is large, even though the probability ratio $\exp(\varepsilon)$ in the LDP definition is high, it remains difficult for an adversary to infer the true location with high probability due to the vast number of possible locations.
Trajectory utility metrics. We follow the Euclidean distance to measure the utility of the perturbed trajectory [37]. The distance between two locations $\tau_{i}$ and $\tau^{\prime}_{i}$ is defined as $\|\tau_{i}-\tau^{\prime}_{i}\|_{2}$. Therefore, given a sensitive trajectory $\mathcal{T}$ and a perturbed trajectory $\mathcal{T}^{\prime}$, the average error (AE) among all locations in the trajectory is:
$AE(\mathcal{T},\mathcal{T}^{\prime})=\frac{1}{|\mathcal{T}|}\sum_{i=1}^{|\mathcal{T}|}\|\tau_{i}-\tau^{\prime}_{i}\|_{2},$
where $|\mathcal{T}|$ is the number of locations in the trajectory. A smaller $AE$ indicates better utility of the perturbed trajectory. Many other trajectory utility metrics, such as Preservation Range Query [9, 37], i.e. whether each perturbed location is within a $\delta$ range of the corresponding sensitive location, are applications of $AE$ and can be derived from it. We defer these results to Appendix C.7 for brevity.
4.1.2 Error Comparison on Synthetic Datasets
(a) $\mathcal{S}=[0,1)\times[0,1)$
(b) $\mathcal{S}=[0,2)\times[0,10)$
Figure 4: Comparison on synthetic datasets.
Figure 5: Condition TraCS-D outperforms TraCS-C.
Figure 6: TraCS-D vs $k$-$R$ + uniform sampling.
Figure 4 presents the error comparison between TraCS and the strawman method. We can observe the error advantage of TraCS over the strawman method as the privacy parameter $\varepsilon$ increases. Specifically, in Figure 4(a), TraCS-C exhibits smaller $AE$ when $\varepsilon\gtrsim 3$, and TraCS-D exhibits smaller $AE$ when $\varepsilon\gtrsim 5$. When the location space expands to $0,2)\times[0,10)$, i.e. Figure 4(b), the advantage of TraCS becomes more significant. Among TraCS, TraCS-C consistently has a smaller $AE$ than TraCS-D. This is because TraCS-D often deals with larger distance spaces $\mathcal{D}_{r(\varphi)}$, which also magnifies the error of direction perturbation. Statistically, the mean $AE$ of TraCS-D is $91.1\%$ of the strawman method in Figure 4(a) and $86.6\%$ in Figure 4(b) across all the $\varepsilon$ values. For TraCS-C the corresponding ratios are $75.5\%$ in Figure 4(a) and $64.0\%$ in Figure 4(b).
Conditions TraCS-D outperforms TraCS-C. The error of TraCS-C depends on the size of the distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, while the error of TraCS-D depends only on the size of the distance space $\mathcal{D}_{r(\varphi)}$. Therefore, if $\mathcal{D}_{r(\varphi)}$ is majorly smaller than $\mathcal{D}_{a}$ or $\mathcal{D}_{b}$, TraCS-D will outperform TraCS-C. We conduct experiments to verify this condition. Specifically, we set $\mathcal{S}=[0,2)\times[0,10)$, and choose a sensitive trajectory along the $x$-axis, i.e. the short side of $\mathcal{S}$. Then we perturb the trajectory $1000$ times using TraCS-C and TraCS-D and calculate the average $AE$. In this case, $|\mathcal{D}_{r(\varphi)}|\approx 2$, much smaller than $|\mathcal{D}_{b}|=10$. Consequently, TraCS-D exhibits smaller $AE$ than TraCS-C across all the $\varepsilon$ values, as shown in Figure 6. In this experiment, the mean $AE$ of TraCS-D is $49.8\%$ of TraCS-C across all the $\varepsilon$ values.
Error of direction perturbation. Figure 6 shows the average error of direction perturbation in TraCS-D compared with the strawman method. We collect the sensitive directions of the trajectories in the location space $\mathcal{S}=[0,1)\times[0,1)$, and perform direction perturbation using TraCS-D and the strawman method. For the strawman method, we use $k$-$R$ with $k=3,6,12$. This experiment reflects empirical error results of mechanism $\mathcal{M}_{\circ}$ in Definition 5 and the $k$-$R$ + uniform sampling. We can observe an exponentially decreasing error of TraCS-D as $\varepsilon$ increases, as stated in Theorem 3. In contrast, the strawman method exhibits an eventually stable error as $\varepsilon$ increases, due to the inherent inner-sector error of $k$-$R$.
(a) TKY dataset
(b) CHI dataset
Figure 7: Comparison on real-world datasets.
4.1.3 Error Comparison on Real-world Datasets
Figure 7 shows the error comparison on the TKY and CHI areas. The trends mirror those observed on the synthetic data. Across all $\varepsilon$ values, the mean $AE$ of TraCS-D is $96.8\%$ of the strawman method on TKY and $94.5\%$ on CHI; for TraCS-C the corresponding ratios are $89.7\%$ and $61.2\%$. The performance gap is smaller on TKY than on CHI, because TKY’s location space is more compact, which lowers the error values for all perturbation methods.
4.1.4 Choose TraCS-D or TraCS-C
From the above experiments, we can conclude criteria for choosing TraCS-D or TraCS-C: (i) TraCS-C generally has better trajectory utility than TraCS-D for random trajectories; (ii) TraCS-D is better for specific trajectories where their distance spaces are smaller than in TraCS-C, and suitable for circular areas.
4.2 Results on Discrete Space
TraCS can be applied to discrete spaces by rounding the perturbed locations to the nearest discrete points. Notably, the privacy guarantee remains valid for the continuous space and hence is also valid for the discrete space, independent of the discretization. We will evaluate the performance of TraCS on discrete spaces and compare it with existing methods.
4.2.1 Setup
Our evaluation includes the following LDP methods for collecting trajectories in discrete spaces:
  • NGram [9]: It is often impractical to acquire reachability knowledge in practice [37]. We instead consider a strong reachability constraint, i.e. from each location, the next location can only be within a distance of $0.75\cdot\hat{a}$, [^4] While the original NGram employs much stricter time-reachability constraints (e.g. the next location must be within 0.10.1h $\cdot$ 44km/h), this constraint is also known to the adversary, which undermines the privacy guarantee. TraCS are LDP mechanisms guaranteeing privacy for the entire domain and do not incorporate such constraints. where $\hat{a}$ is the maximal length of the location space. This constraint significantly reduces the location space for perturbation, especially for the locations near the boundary.
  • L-SRR [29]: We use group number $g=3$, the same as the empirical optimal value in their paper and code. Locations in the first group are within $0.3\cdot\hat{a}$ distance from the sensitive location, where $\hat{a}$ is the maximal length of the location space. The second group is within $0.6\cdot\hat{a}$ distance from the sensitive location, and the remaining locations are in the third group.
  • ATP [37]: We set the $k$-$R$ parameter $k=6$ in ATP. The privacy parameter $\varepsilon$ in original ATP is for the entire trajectory, we distribute it to each location’s perturbation according to their algorithms.
  • TraCS: We use the same privacy parameter $\varepsilon_{d}$ as in the continuous space experiments. For each perturbed location from TraCS, we round it to the nearest discrete location.
For a fair comparison, we use $\varepsilon$ as the privacy parameter for each location in all methods. The comparison is conducted on the following datasets:
  • Synthetic datasets: We uniformly discretize $0,1)\times[0,1)$ into $m$ discrete points and treat the $m$ points as the discrete location space. Specifically, we set $m=10\times 10$ and $m=60\times 60$ for comparison. Along with them, we generate $100$ random trajectories with $100$ locations each as synthetic datasets.
  • TKY and CHI: We treat every location in the TKY and CHI datasets as a discrete point, which together constitute the discrete location space. TKY contains $7,798$ discrete locations; CHI contains $1,000$ discrete locations. We use the first $100$ trajectories for evaluation.
For the TKY and CHI datasets, we perform the perturbation $5$ times for each trajectory and compute the average $AE$, which is the same setup as in ATP [37].
4.2.2 Error Comparison on Synthetic Datasets
(a) $m=10\times 10$
(b) $m=60\times 60$
Figure 8: Comparison on synthetic datasets. $m$ is the number of discrete points in the location space.
Figure 8 presents the error comparison on the synthetic dataset. Figure 8(a) shows results for $m=10\times 10$, and Figure 8(b) shows results for $m=60\times 60$. We observe that TraCS-C consistently outperforms NGram, L-SRR, and ATP across all the $\varepsilon$ values, and TraCS-D also outperforms them when $\varepsilon\gtrsim 3$. Moreover, the performance advantage of TraCS becomes more significant as $m$ increases to $60\times 60$. This is because finer-grained discretization allows TraCS’s rounding to align with the sensitive locations more accurately. Statistically, the mean $AE$ of TraCS-C is $66.1\%$ of NGram, $57.1\%$ of L-SRR, and $63.8\%$ of ATP in Figure 8(a), and the proportions are $57.6\%$, $52.6\%$, and $61.1\%$ in Figure 8(b). The mean $AE$ of TraCS-D is $76.9\%$ of NGram, $66.4\%$ of L-SRR, and $74.2\%$ of ATP in Figure 8(a), and the proportions are $71.2\%$, $65.0\%$, and $75.5\%$ in Figure 8(b). These results indicate the advantage of TraCS for synthetic trajectories in discrete spaces.
4.2.3 Error Comparison on Real-world Datasets
Figure 9 presents the error comparison on the TKY and CHI datasets. Compared to the synthetic dataset, NGram, L-SRR, and ATP achieve lower errors, particularly when $\varepsilon$ is small. However, as $\varepsilon$ increases, TraCS still outperforms them.
Reasons for the results. (i) ATP performs exceptionally well when $\varepsilon$ is small, particularly on the TKY dataset. This is because this dataset contains many “condensed” locations, i.e. no other locations near the trajectory. In such cases, if the bi-direction perturbation of ATP is accurate, the perturbed location is error-free. (ii) The $AE$ of ATP does not decrease significantly as $\varepsilon$ increases, which is consistent with their paper. NGram shows a similar trend, especially on the TKY dataset. This is due to the low performance of the Exponential mechanism for large location spaces, particularly when the locations are concentrated in a small area. In such cases, the score function struggles to differentiate the scores among the locations. L-SRR exhibits a similar trend. (iii) TraCS exhibits larger $AE$ than them when $\varepsilon$ is small because it focuses on the rectangular area formed by the longitude and latitude, instead of a set of discrete locations in the cities. Therefore, the perturbed locations may be far from the city center (where locations are condensed and benefits discrete mechanisms) when $\varepsilon$ is small.
(a) TKY dataset
(b) CHI dataset
Figure 9: Comparison on real-world datasets. TraCS operates on rectangular areas encompassing the city, which may result in larger AEAE when $\varepsilon$ is small.
Table 2: Time cost comparison (in milliseconds).
ATP NGram L-SRR TraCS-D TraCS-C
Total $145.7$ $100.9$ $6.2$ $0.06$ $0.05$
Perturb $125.8$ $92.8$ $0.003$ $0.018$ $0.003$
4.2.4 Time Cost Comparison
Table 2 presents the time cost comparison on the TKY dataset. The time cost is measured in milliseconds and averaged per location. It shows that TraCS-D and TraCS-C are significantly faster than NGram, ATP, and L-SRR. Compared to NGram and ATP, this efficiency is primarily due to the piecewise-based mechanisms in TraCS, which require negligible time for perturbation compared to the Exponential mechanism. Compared to L-SRR, TraCS-D and TraCS-C are also faster because they do not require grouping locations.
ATP has the highest time cost because it needs to process two copied trajectories and merges them. NGram has a lower time cost than ATP but is still significantly higher than TraCS. The perturbation procedure accounts for $86.3\%$ and $92.0\%$ of the total time cost in ATP and NGram, respectively. L-SRR has the lowest time cost among the three state-of-the-art methods due to its efficient sampling mechanism, however, its grouping procedure still incurs a non-negligible time cost.
The total time cost of TraCS is less than $0.05\%$ of ATP and NGram, and less than $1\%$ of L-SRR. The perturbation procedure of TraCS-D and TraCS-C only takes $30\%$ and $6\%$ time, respectively. TraCS-D has a higher time cost than TraCS-C because it needs to calculate the distance space $\mathcal{D}_{r(\varphi)}$ for each location in the trajectory, while TraCS-C can use the same $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ for all locations.
4.3 Summary of Evaluation
From the evaluation results on both continuous and discrete spaces, we can summarize the following findings:
  • In continuous spaces, TraCS generally exhibits better trajectory utility than the strawman method, particularly when the privacy parameter $\varepsilon$ is large. Meanwhile, TraCS-C has better utility than TraCS-D for common-shape location space and trajectories.
  • In discrete spaces, TraCS generally outperforms NGram, L-SRR, and ATP, particularly when the privacy parameter $\varepsilon$ is large. Furthermore, TraCS has significantly lower time cost, requiring less than $1\%$ of their time cost.
5 Conclusions
This paper presents TraCS-D and TraCS-C, providing local differential privacy guarantees for trajectory collection in continuous location spaces. TraCS-D uses a novel direction-distance perturbation procedure, while TraCS-C perturbs the Cartesian coordinates of each location. These two methods can also be applied to discrete location spaces by rounding each perturbed location to the nearest discrete point. In this case, they are more efficient than existing discrete methods. Trajectory utility of TraCS is theoretically analyzed and empirically evaluated. Evaluation results on discrete location spaces show that TraCS outperforms state-of-the-art methods in trajectory utility, particularly when the privacy parameter is large.
Ethical Considerations
This paper does not involve any human subjects or animals. While we use real-world trajectory data, they are publicly available, and all personal information has been anonymized.
Open Science
All code for the trajectory collection methods, evaluations, and datasets used in this paper will be made publicly available upon publication. We will include scripts for reproducing the experiments and generating the figures presented in the paper. An anonymized snapshot for review is currently available at https://anonymous.4open.science/r/TraCS-USENIX/README.md.
Appendix A Related Work
This section reviews related trajectory collection methods under LDP and other quantifiable privacy notions. More privacy-preserving trajectory collection methods can be found in recent surveys [26, 3].
Trajectory Collection under LDP
We have disscussed three state-of-the-art trajectory collection methods under pure LDP: NGram [9], L-SRR [29], and ATP [37]. Among them, L-SRR was originally applied for origin-destination location pairs, while NGram, ATP, and this paper focus on trajectory-level perturbation. NGram attempts to ensure pure LDP for the “trajectory space”, also named as trajectory-level or instance-level LDP in surveys [26, 3], which is a stronger privacy than the location space. However, this is proved extremely challenging as the trajectory space increases exponentially with the number of possible locations in the location space. L-SRR, ATP and this paper focus on ensuring pure LDP for the location space. Unlike L-SRR, ATP and other existing methods based on discrete mechanisms, this paper focuses on continuous location spaces, which are stronger than any fine-grained discrete location space.
Another line of research is trajectory synthesis for privacy-preserving trajectory publication. A typical work is LDPTrace [10], which synthesizes trajectories using a generative model. LDPTrace discretizes the location space into cells and represents each trajectory as a sequence of cells. Transition patterns between neighboring cells are perturbed using LDP mechanisms and collected by an untrusted curator. The curator then trains a Markov chain model on the perturbed transition patterns to capture movement behaviors. Synthetic trajectories are generated from the trained model for publication. However, compared to methods that perturb each trajectory, such as NGram, L-SRR, and ATP, synthetic trajectories may be overly too random and not reflect the original trajectories.
Trajectory Collection under DP
As a weaker privacy notion than LDP, (central) DP [12, 13] assumes a trusted curator who processes the raw data and releases the perturbed data. It requires that only neighboring locations in the location space be indistinguishable, rather than any two arbitrary locations. Compared to LDP, the concept of “neighboring” in DP has many interpretations. A comprehensive survey on DP-based trajectory collection and publication is available in [26].
The usage of direction information in trajectory collection dates back to an early work SDD [17], which attempts to add unbounded Laplace noise to the direction information. It provides a DP guarantee for unbounded continuous location spaces, but results in poor utility. The notion of “n-gram” in trajectory collection dates back to [5], where it abstracts a trajectory pattern as a gram. The idea of hierarchical decomposition can also be found in [36], where a spatial decomposition method is used for trajectory synthesis.
External knowledge is also widely used in DP-based trajectory collection methods. NGram’s prior work [8] provides a DP-based solution for trajectory synthesis by exploiting the public knowledge of the road network. Focusing on particular privacy-concerned locations, or sensitive zones [30], can also improve the trajectory utility while maintaining partial privacy guarantees.
Trajectory Collection under Other Privacy Notions
A widely used relaxed privacy notion than LDP for trajectory collection is Geo-indistinguishability [2, 25], where the indistinguishability between two locations is a function of their distance, i.e. privacy can decrease with distance. Input-discriminative LDP [14] goes one step further, assigning each location its own privacy level. Threshold-integrated LDP [35] defines a threshold to constrain the sensitive region near a location. This is similar to NGram’s reachability constraint but with a formal privacy definition. More relaxed LDP notions for trajectory collection can be found in paper survey [26]. Although these relaxed LDP notions generally provide better trajectory utility than pure LDP, they sacrifice the strong privacy guarantee of pure LDP, and often need hyperparameters to define the privacy level.
Appendix B Proofs
Table 3: Notations used in this paper.
Notation Description
$\mathcal{T}\coloneqq\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$ Sensitive trajectory
$\mathcal{T}^{\prime}\coloneqq\{\tau_{1}^{\prime},\tau_{2}^{\prime},\ldots,\tau_{n}^{\prime}\}$ Perturbed trajectory
$\mathcal{S}\coloneqq[a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$ Continuous location space
$\mathcal{M}_{\circ}$ Direction perturbation mechanism
$\mathcal{M}_{-}$ Distance perturbation mechanism
$\varphi\in\mathcal{D}_{\varphi}\coloneqq[0,2\pi)$ Direction & direction space
$\varphi^{\prime}$ Perturbed direction
$r(\varphi)\in\mathcal{D}_{r(\varphi)}$ Distance & distance space along $\varphi$
$r^{\prime}(\varphi)$ Perturbed distance along $\varphi$
$\overline{r}(\varphi)\in[0,1)$ Normalized distance over $\mathcal{D}_{r(\varphi)}$
Table 3 summarizes the notations used throughout the remainder of the Appendix.
B.1 LDP Proof of Definition 5
The proof applies to all piecewise-based mechanisms, including the direction perturbation mechanism in Definition 5 and the distance perturbation mechanism in Definition 6.
Proof.
For any two sensitive directions $\varphi_{1}$, $\varphi_{2}$ and any $\varphi^{\prime}\in[0,2\pi)$, we have:
$\frac{pdf[\mathcal{M}_{\circ}(\varphi_{1})=\varphi^{\prime}]}{pdf[\mathcal{M}_{\circ}(\varphi_{2})=\varphi^{\prime}]}\leq\frac{p_{\varepsilon}}{p_{\varepsilon}/\exp(\varepsilon)}=\exp(\varepsilon).$
The inequality holds because any $\varphi^{\prime}$ is sampled with a probability of at most $p_{\varepsilon}$ and at least $p_{\varepsilon}/\exp(\varepsilon)$, regardless of whether the input is $\varphi_{1}$, $\varphi_{2}$ or any other value. ∎
B.2 Proof of Theorem 2
Proof.
Denote TraCS-D in Algorithm 1 as $\mathcal{M}$. We need to prove that $\mathcal{M}$ satisfies $\varepsilon$-LDP for each location $\tau\in\mathcal{S}$, i.e.
$\forall\tau_{1},\tau_{2},\tau^{\prime}\in\mathcal{S}:\frac{pdf[\mathcal{M}(\tau_{1})=\tau^{\prime}]}{pdf[\mathcal{M}(\tau_{2})=\tau^{\prime}]}\leq\exp(\varepsilon).$
In TraCS-D, each location $\tau$ is represented in a direction-distance space with unique coordinates $(\varphi,r(\varphi))$. The key insight of the proof is that the perturbation of $\varphi$ and $r(\varphi)$ satisfies $\varepsilon_{d}$-LDP and $(\varepsilon-\varepsilon_{d})$-LDP, respectively. Then TraCS-D ensures $\varepsilon$-LDP for the location space $\mathcal{S}=\mathcal{D}_{\varphi}\otimes\mathcal{D}_{r(\varphi)}$. We provide proofs by the Composition Theorem 1 and by computing the 2D pdfpdf ratios in the LDP definition.
By Composition Theorem 1. The definitions of piecewise-based mechanisms $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ imply that they satisfy LDP. Specifically, for $\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$,
$\forall\varphi_{1},\varphi_{2},\varphi^{\prime}\in[0,2\pi):\frac{pdf[\mathcal{M}_{\circ}(\varphi_{1})=\varphi^{\prime}]}{pdf[\mathcal{M}_{\circ}(\varphi_{2})=\varphi^{\prime}]}\leq\exp(\varepsilon_{d}),$
is guaranteed by Definition 5 of $\mathcal{M}_{\circ}$. Similarly, $\mathcal{M}_{-}(\overline{r}(\varphi);\varepsilon-\varepsilon_{d})$ in Definition 6 ensures $(\varepsilon-\varepsilon_{d})$-LDP for $\overline{r}(\varphi)$. For clarity, let $\overline{r}\coloneqq\overline{r}(\varphi)$, which yields
$\forall\overline{r}_{1},\overline{r}_{2},\overline{r}^{\prime}\in[0,1):\frac{pdf[\mathcal{M}_{-}(\overline{r}_{1})=\overline{r}^{\prime}]}{pdf[\mathcal{M}_{-}(\overline{r}_{2})=\overline{r}^{\prime}]}\leq\exp(\varepsilon-\varepsilon_{d}).$
The subsequent linear mapping from $\overline{r}\in[0,1)$ to $r\in\mathcal{D}_{r(\varphi^{\prime})}$ is a post-processing step that does not affect the LDP property. Specifically, denote the linear mapping as $g:[0,1)\to\mathcal{D}_{r(\varphi^{\prime})}$, and denote $r_{1}=g(\mathcal{M}_{-}(\overline{r}_{1}))$ and $r_{2}=g(\mathcal{M}_{-}(\overline{r}_{2}))$ as two random variables. It is clear that
$\forall r_{1},r_{2},r^{\prime}\in\mathcal{D}_{r(\varphi^{\prime})}:\frac{pdf[r_{1}=r^{\prime}]}{pdf[r_{2}=r^{\prime}]}\leq\exp(\varepsilon-\varepsilon_{d}),$
because the linear mapping $g$ is a deterministic function without randomness. Thus, $\mathcal{M}_{-}$ ensures $(\varepsilon-\varepsilon_{d})$-LDP for $\mathcal{D}_{r(\varphi^{\prime})}$.
Combining these results with Sequential Composition Theorem 1, we can conclude that $\mathcal{M}=(\mathcal{M}_{\circ},\mathcal{M}_{-})$ ensures $\varepsilon$-LDP for each location $\tau=(\varphi,r(\varphi))\in\mathcal{S}$. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP.
By computing the 2D pdfpdf ratio. Assuming two sensitive locations $\tau_{1}=(\varphi_{1},\overline{r}_{1})$ and $\tau_{2}=(\varphi_{2},\overline{r}_{2})$, where $\overline{r}_{1}$ and $\overline{r}_{2}$ are normalized distances from the reference location along $\varphi_{1}$ and $\varphi_{2}$, respectively. For any output $\tau^{\prime}=(\varphi^{\prime},\overline{r}^{\prime})$, the 2D pdfpdf of getting it from $\tau_{1}$ and $\tau_{2}$ are
$\begin{split}pdf[\mathcal{M}(\tau_{1})=\tau^{\prime}]=pdf[\mathcal{M}_{\circ}(\varphi_{1})=\varphi^{\prime}]\cdot pdf[\mathcal{M}_{-}(\overline{r}_{1})=\overline{r}^{\prime}],\\ pdf[\mathcal{M}(\tau_{2})=\tau^{\prime}]=pdf[\mathcal{M}_{\circ}(\varphi_{2})=\varphi^{\prime}]\cdot pdf[\mathcal{M}_{-}(\overline{r}_{2})=\overline{r}^{\prime}],\end{split}$
where $\mathcal{M}_{-}$ is applied to the same normalized distance space (from the reference location to the boundary of $S$ along $\varphi^{\prime}$). The ratio is bounded by $\exp(\varepsilon_{d})\cdot\exp(\varepsilon-\varepsilon_{d})=\exp(\varepsilon)$ due to the definition of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
B.3 Proof of Theorem 3
Proof.
We prove the $\Theta(e^{-\varepsilon/2})$ convergence rate by calculating the MSE of the perturbation mechanism. Specifically, the MSE of $\mathcal{M}_{-}$ is calculated as follows:
$\mathrm{MSE}[\mathcal{M}_{-}(\overline{r})]=\int_{0}^{1}(\overline{r}^{*}-\overline{r})^{2}pdf[\mathcal{M}_{-}(\overline{r})=\overline{r}^{*}]d\overline{r}^{*}.$
The worst-case MSE is achieved when $\overline{r}=0$ or $\overline{r}=1$, i.e. at the endpoints (similar to the variance calculation in [31]). In this case,
$\begin{split}\mathrm{MSE}[\mathcal{M}_{-}(0)]&=\int_{0}^{2C}\left(\overline{r}^{*}-0\right)^{2}p_{\varepsilon}d\overline{r}^{*}+\int_{2C}^{1}\left(\overline{r}^{*}-0\right)^{2}\frac{p_{\varepsilon}}{e^{\varepsilon}}d\overline{r}^{*}\\ &=\frac{8C^{3}p_{\varepsilon}}{3}+\frac{(1-8C^{3})p_{\varepsilon}}{3e^{\varepsilon}}.\end{split}$
with $p_{\varepsilon}$ and $C$ defined in Definition 6. We omit the calculation for $\overline{r}=1$ as it is symmetric to $\overline{r}=0$. Plugging in the values of $p_{\varepsilon}$ and $C$, we can simplify the MSE using big $\mathcal{O}$ notation:
$\begin{split}\mathrm{MSE}[\mathcal{M}_{-}(0)]&=\frac{1}{3}e^{-\varepsilon/2}+\frac{(e^{\varepsilon/2}-1)^{3}}{3e^{\varepsilon/2}(e^{\varepsilon}-1)^{2}}\\ &=\frac{1}{3}e^{-\varepsilon/2}+\mathcal{O}(e^{-\varepsilon}).\end{split}$
Therefore, as $\varepsilon\to\infty$, the worst-case MSE of $\mathcal{M}_{-}$ converges to zero at a rate of $\Theta(e^{-\varepsilon/2})$. The MSE of $\mathcal{M}_{\circ}$ can be calculated in the same way, resulting in a convergence rate of $\Theta(e^{-\varepsilon/2})$ as well. ∎
B.4 Proof of Theorem 4
Proof.
Denote TraCS-C in Algorithm 2 as $\mathcal{M}$. We need to prove that $\mathcal{M}$ satisfies $\varepsilon$-LDP for each location $\tau\in\mathcal{S}$. In TraCS-C, each location $\tau$ is represented in Cartesian space $\mathcal{D}_{a}\times\mathcal{D}_{b}$ with unique coordinates $(a,b)$. Therefore, the key insight is to prove that the perturbation of $a$ and $b$ satisfies $\varepsilon/2$-LDP each.
In Theorem 2, we have already proved that $\mathcal{M}_{-}(\overline{a},\varepsilon/2)$ satisfies $\varepsilon/2$-LDP for the normalized distance $\overline{a}$. Then, the linear mapping from $\overline{a}\in[0,1)$ to $a\in\mathcal{D}_{a}$ is a post-processing step that does not affect the LDP property. This means $\mathcal{M}_{-}$ ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{a}$. In the same way, it also ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{b}$. Therefore, $\mathcal{M}=(\mathcal{M}_{-},\mathcal{M}_{-})$ satisfies $\varepsilon$-LDP for each location $\tau=(a,b)\in\mathcal{S}$ by Sequential Composition Theorem 1. (Alternatively, it can also be proved by directly computing the 2D pdfpdf ratio for obtaining the same location via $\mathcal{M}=(\mathcal{M}_{-},\mathcal{M}_{-})$.) Therefore, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
Appendix C Complementary Materials
(a) $m=4\times 4$ (gray points)
(b) $m=11\times 11$ (gray points)
Figure 10: The Exponential mechanism $\mathcal{M}_{\text{exp}}(0,0)$ with $\varepsilon=4$ on two discrete location spaces within $0,1]\times[0,1]$. The number and arrangement of points affect the mechanism.
(a) $\mathcal{S}=[0,1]\times[0,1]$
(b) $\mathcal{S}=[0,1]\times[0,1]$
Figure 11: TraCS-C mechanism $\mathcal{M}_{\text{TraCS-C}}(0,0)$ with $\varepsilon=4$ on $\mathcal{S}=[0,1]\times[0,1]$, which encompasses both discrete location spaces in Figure 10. The sampling probability is defined over areas rather than individual points.
C.1 Weaknesses of the Exponential Mechanism (Section 2.3)
C.1.1 High Computational Complexity
The biggest weakness of the Exponential mechanism is the complexity in calculating the score function and sampling from the probability distribution. Before perturbing location $x$, the Exponential mechanism requires computing the score function $d(x,y)$ for every possible location $y$ in the location space. If the location space has $m$ locations, the time complexity of computing the score function for a single location is $\mathcal{O}(m)$, which is computationally expensive for large-scale spaces. Moreover, ATP’s dynamic strategy to constrain the location space for perturbation makes it need to recompute the score function for each location, resulting in $\mathcal{O}(m)$ time complexity, where $m$ is the size of the location space.
The complexity of sampling from the Exponential mechanism is also $\mathcal{O}(m)$, as its cumulative distribution function (CDF) is an $m$-piece function. Sampling requires comparing a random number with the CDF values of all $m$ locations.
In contrast, we will see that piecewise-based perturbation mechanisms in TraCS do not require score functions, and they sample from a $3$-piece CDF at each perturbation, resulting in a negligible constant time complexity, i.e. $\Theta(1)$, for each perturbation.
C.1.2 Affected by the Number and Arrangement of Locations
Because the sampling probability depends on the score function, the spatial distribution and relative distances between locations further affect the probability assigned to each location.
Figure 10 illustrates the Exponential mechanism $\mathcal{M}_{\text{exp}}$ applied to the sensitive location $\tau=(0,0)$ with $\varepsilon=4$ on two discrete location spaces within $0,1]\times[0,1]$. Specifically, Figure 10(a) presents a $4\times 4$ grid of locations, while Figure 10(b) depicts an $11\times 11$ grid. We can observe that the sampling probabilities of $\mathcal{M}_{\text{exp}}$ differ for the same sensitive location $\tau=(0,0)$, even though both location spaces are contained within the same $0,1]\times[0,1]$ area. When the number of locations increases and the arrangement becomes denser (i.e. $m=11\times 11$), the sampling probabilities for each location become smaller and more uniform.
In contrast, Figure 11 illustrates the TraCS-C mechanism $\mathcal{M}_{\text{TraCS-C}}$ applied to the same sensitive location $\tau=(0,0)$ with $\varepsilon=4$ over the $0,1]\times[0,1]$ area. Unlike the Exponential mechanism, TraCS-C is designed for continuous spaces, where the sampling probability is defined over areas rather than individual points.
To apply TraCS-C to the discrete spaces shown in Figure 10, one can simply round each perturbed location to its nearest discrete point. Since this rounding is a post-processing step, it does not affect the LDP guarantee.
C.2 Dominant Sector Comparison Between the Strawman Approach and TraCS-D (Example 1)
(a) $k=6$
(b) $k=12$
Figure 12: Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach. TraCS-D (solid line) achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach (dashed line), having better adaptiveness across different $\varepsilon$ regions.
For direction perturbation, a small-size and high-probability dominant sector is desirable for more accurate perturbation. However, these two properties are generally conflicting with each other due to the LDP constraint: a smaller dominant sector typically has a lower probability. We will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value. Figure 12 shows two exemplary quantitative comparisons between them.
Small $\varepsilon$ region. In this case, the dominant sector of TraCS-D (red solid line) is larger than the strawman approach, but the probability of the dominant sector (black solid line) is significantly higher, especially when $k$ is large. This indicates that, although the dominant sector of TraCS-D is larger in this region, it is sampled with a higher probability.
Large $\varepsilon$ region. In such cases, the dominant sector of TraCS-D is significantly smaller than the strawman approach, while maintaining almost the same probability as the strawman approach.
We can observe the effect of $k$ on the strawman approach. Larger $k$ values in the strawman approach reduce the inner-sector error compared to TraCS-D, but also significantly decrease the probability of the dominant sector when $\varepsilon$ is small. Conversely, smaller $k$ values increase the probability of the dominant sector when $\varepsilon$ is small, but result in a large inner-sector error when $\varepsilon$ is large. In contrast, TraCS-D adaptively balances these trade-offs by adjusting the dominant sector’s size and probability according to $\varepsilon$.
C.3 Detailed Form of $|\mathcal{D}_{r(\varphi)}|$ (Section 3.2.2)
For each location $\tau_{i}$, denote its directions to the four endpoints of the rectangular location space as $\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4}$. We have:
$\begin{split}\varphi_{1}&=\mathrm{atan2}\left(b_{\text{end}}-b_{i},a_{\text{end}}-a_{i}\right),\\ \varphi_{2}&=\mathrm{atan2}\left(b_{\text{end}}-b_{i},a_{\text{sta}}-a_{i}\right),\\ \varphi_{3}&=\mathrm{atan2}\left(b_{\text{sta}}-b_{i},a_{\text{sta}}-a_{i}\right)+2\pi,\\ \varphi_{4}&=\mathrm{atan2}\left(b_{\text{sta}}-b_{i},a_{\text{end}}-a_{i}\right)+2\pi.\end{split}$
Then, $|\mathcal{D}_{r(\varphi)}|$ can be derived using trigonometric functions, which leads to the following four cases:
$\begin{split}|\mathcal{D}_{r(\varphi)}|=\begin{dcases}\frac{a_{\text{end}}-a_{i}}{\cos\varphi}&\text{if}\ \varphi\in[0,\varphi_{1})\cup[\varphi_{4},2\pi),\\ \frac{b_{\text{end}}-b_{i}}{\sin\varphi}&\text{if}\ \varphi\in[\varphi_{1},\varphi_{2}),\\ \frac{a_{i}-a_{\text{sta}}}{-\cos\varphi}&\text{if}\ \varphi\in[\varphi_{2},\varphi_{3}),\\ \frac{b_{i}-b_{\text{sta}}}{-\sin\varphi}&\text{if}\ \varphi\in[\varphi_{3},\varphi_{4}),\end{dcases}\end{split}$
depending on the direction $\varphi$ from reference location $\tau_{i}=(a_{i},b_{i})$ to the boundary of the rectangular location space $a_{\text{sta}},a_{\text{end}}]\times[b_{\text{sta}},b_{\text{end}}]$.
C.4 More Extensions (Section 3.5.5)
C.4.1 Other Piecewise-based Mechanisms
In Definition 5 and Definition 6, we use specific instantiations for the parameters $p$ and $[l,r)$ in the piecewise-based mechanism. Besides these instantiations, TraCS can be integrated with any other piecewise-based mechanisms [31, 21, 24, 23, 38] by redesigning them. For example, SW [21] is a famous mechanism proposed for distribution estimation; it is defined on $0,1)\to[-b,1+b)$, with $b$ as a parameter. It uses different $p$ and $[l,r)$, and employs a strategy of maximizing the mutual information between the perturbed and sensitive data to determine $b$. We can redesign SW for the direction perturbation and distance perturbation in TraCS-D and TraCS-C. Appendix C.5 provides the detailed redesign of SW for them and shows intuitive comparisons with Definition 5 and Definition 6. The experimental section also provides an comparison between the redesigned SW and TraCS-D.
C.4.2 Other Shapes of Location Space
The design of LDP mechanisms is space-specific, implying that extensions to other shapes of location space are non-trivial.
For TraCS, the challenge lies in determining the size of each subspace, i.e. $|\mathcal{D}_{\varphi}|$, $|\mathcal{D}_{r(\varphi)}|$, $|\mathcal{D}_{a}|$, and $|\mathcal{D}_{b}|$. Among them, $|\mathcal{D}_{r(\varphi)}|$ is not apparent but can be calculated for rectangular location spaces. Although calculating $|\mathcal{D}_{r(\varphi)}|$ is generally non-trivial for irregular shapes, TraCS can be extended to parallelogram location spaces by a linear transformation to a rectangular space. Meanwhile, TraCS-D can be used for circular areas, which are widely used in relative location representation, e.g. radar and sonar systems.
TraCS-D for circular area. When the location space is circular with radius $R$, TraCS-D has a more straightforward design. The main insight is that such a location space $\mathcal{S}_{R}$ has a natural representation in polar coordinates, i.e. $\mathcal{S}_{R}=\{(\varphi,r):\varphi\in[0,2\pi),r\in[0,R)\}$. Any location $\tau_{i}\in\mathcal{S}_{R}$ corresponds to a pair of coordinates $(\varphi_{i},r_{i})$, and this representation is unique and reversible. In this case, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ and the distance space $\mathcal{D}_{r}=[0,R)$ are independent, so we can apply $\mathcal{M}_{\circ}$ to $\varphi_{i}$ and $\mathcal{M}_{-}$ to $r_{i}/R$. The perturbed location $\tau_{i}^{\prime}$ is then $(\varphi_{i}^{\prime},r_{i}^{\prime})=(\mathcal{M}_{\circ}(\varphi_{i}),\mathcal{M}_{-}(r_{i}/R)\cdot R)$, still within polar coordinates.
C.5 Redesigned SW (Appendix C.4.1)
We redesign the SW mechanism to achieve $\varepsilon$-LDP for the circular space $0,2\pi)$ and the linear space $[0,1)$. The key idea is to transform the original sampling distribution to these spaces while preserving the LDP constraint. Similar to Definition 5, the redesigned SW mechanism for direction perturbation is defined as follows.
Definition 7 (Redesigned SW for direction perturbation).
Given a sensitive direction $\varphi$ and a privacy parameter $\varepsilon$, redesigned SW for direction perturbation is a mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ defined by:
$pdf[\mathcal{M}_{\circ}(\varphi)=\varphi^{\prime}]=\begin{dcases}p_{\varepsilon}&\text{if}\ \varphi^{\prime}\in[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon}),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=\frac{1}{2\pi\varepsilon}(\exp(\varepsilon)-1)$ is the sampling probability, and $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
$\begin{split}l_{\varphi,\varepsilon}&=\left(\varphi-\pi\frac{\exp(\varepsilon)(\varepsilon-1)+1}{(\exp(\varepsilon)-1)^{2}}\right)\mod 2\pi,\\ r_{\varphi,\varepsilon}&=\left(\varphi+\pi\frac{\exp(\varepsilon)(\varepsilon-1)+1}{(\exp(\varepsilon)-1)^{2}}\right)\mod 2\pi.\end{split}$
Similar to Definition 6, the redesigned SW mechanism for distance perturbation is defined as follows.
Definition 8 (Redesigned SW for distance perturbation).
Given a sensitive distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, redesigned SW for distance perturbation is a mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ that
$pdf[\mathcal{M}_{-}(\overline{r}(\varphi))=\overline{r}^{\prime}(\varphi)]=\begin{dcases}p_{\varepsilon}&\text{if}\ \overline{r}^{\prime}(\varphi)\in[u,v),\\ p_{\varepsilon}/\exp{(\varepsilon)}&\text{otherwise},\end{dcases}$
where $p_{\varepsilon}=(\exp(\varepsilon)-1)/\varepsilon$ is the sampling probability, and $[u,v)$ is the sampling interval that
$\begin{split}[u,v)&=\begin{cases}\overline{r}(\varphi)+[-C,C)&\text{if}\ \overline{r}^{\prime}(\varphi)\in[C,1-C),\cr[0,2C)&\text{if}\ \overline{r}^{\prime}(\varphi)\in[0,C),\cr[1-2C,1)&\text{otherwise},\end{cases}\end{split}$
with $C=(\exp(\varepsilon)(\varepsilon-1)+1)/(2(\exp(\varepsilon)-1)^{2})$.
Compared with Definition 5 and Definition 6, the redesigned SW mechanisms have a larger $p_{\varepsilon}$ and a narrower $[l,r)$ or $[u,v]$ interval for sampling. We can treat the redesigned SW mechanisms as a more aggressive perturbation mechanism: it tries to sample from a narrow dominant sector with a higher probability ($p_{\varepsilon}$), which also leads to a higher probability ($p_{\varepsilon}/\exp(\varepsilon)$) of non-dominant sectors due to the LDP constraint.
C.6 Experimental Results for Extensions (Appendix C.4)
C.6.1 Effect of Different Piecewise-based Mechanisms
Figure 14 compares the performance of different piecewise-based mechanisms for TraCS-D. In addition to the mechanisms defined in Definition 5 and Definition 6, we name the redesigned SW for TraCS-D as TraCS-D (R-SW), as detailed in Appendix C.5. We can observe that TraCS-D consistently exhibits smaller errors than TraCS-D (R-SW) across all the $\varepsilon$ values. Statistically, the mean $AE$ of TraCS-D is $52.6\%$ of TraCS-D (R-SW) across all the $\varepsilon$ values. This difference comes from the MSE of their sampling distributions, where Definition 5 and Definition 6 have smaller MSE than R-SW.
C.6.2 TraCS-D for Circular Area
Figure 14 illustrates an example of TraCS-D applied to a circular area $\mathcal{S}=[0,2\pi)\otimes[0,1)$. We set the sensitive location $\tau$ as $(\pi,0.5)$ and the privacy parameter $\varepsilon=5$, then collect $50$ random samples of perturbed locations using TraCS-D. In this case, the dominant area is $0.87\pi,1.13\pi)\otimes[0.32,0.67)$, as determined by the perturbation mechanisms in Definition 5 and Definition 6. We can observe that the perturbed locations are concentrated in the dominant area.
Figure 13: Comparison between different piecewise-based mechanisms.
Figure 14: Fifty samples of TraCS-D at $\tau=(\pi,0.5)$ for circular area.
C.7 Experimental Results for Other Metrics (Section 4.1.1)
A smaller $AE$ generally indicates better preservation of range queries and hotspots [9, 37]. (i) For range query preservation (Formula 17 in [9]), given a threshold $\delta$, a perturbed location is considered correct if it lies within a $\delta$ distance of the corresponding sensitive location. Thus, when $AE\leq\delta$, the perturbed location is expected to satisfy the range query, and a smaller $AE$ implies higher correctness. (ii) For hotspot preservation (Section 6.2.4 in [37]), which measures how many hotspots from a given set remain after perturbation, a smaller $AE$ leads to better hotspot retention after rounding in TraCS.
(a) Range query preservation
(b) Hotspot preservation
Figure 15: Comparison of other trajectory metrics on the CHI dataset. Higher range query preservation and lower error in hotspot preservation indicate better performance.
We further evaluate the performance of TraCS-D and existing methods under varying privacy parameters $\varepsilon$ using these two metrics. Specifically, we use the CHI dataset and adopt the following experimental setups:
  • Range query preservation: We set the threshold $\delta=0.1$, meaning a perturbed location is considered correct if its distance from the sensitive location is less than $0.1$ in longitude-latitude coordinates.
  • Hotspot preservation: The hotspot set is defined as the first $50\%$ of locations in the CHI dataset. We count the number of hotspots before and after perturbation for the first $100$ trajectories and report the difference.
The results are shown in Figure 15. It can be seen that TraCS generally outperforms existing methods across different privacy levels, demonstrating superior effectiveness in preserving both range queries and hotspots. Particularly, TraCS shows significant improvement in hotspot preservation compared to existing methods across all privacy levels.